### v4

# Tony Lindeberg

Abstract
When a visual observer views objects and spatio-temporal events in the environment, the image data may undergo substantial variabilities, as caused by variations in the viewing distance, the viewing direction and the relative motion between the object or event and the observer. A non-trivial question concerns how the vision system should be able to establish an identity between the receptive field responses obtained from different observations of the same object or event under such variations in the viewing conditions, since the responses of the receptive fields in the first layers of the visual hierarchy will be strongly influenced by the underlying geometric image transformations.
This paper presents a systematic methodology for addressing this problem, by giving an overview of a principled theory for modelling the interaction between geometric image transformations and receptive field responses, for a visual observer that views objects and spatio-temporal events in the environment. Specifically, the paper presents fundamental results regarding the influence on the receptive field responses due to 4 main classes of geometric image transformations, which represent the main sources of variabilities in the geometry between the object or the event and the observer, based on local approximations of the perspective or projective image projection models between different views.
By postulating that the family of receptive fields should be well-behaved (covariant) under these classes of geometric image transformations, it follows that the receptive field shapes should be expanded over the degrees of freedom of the corresponding image transformations, to enable a formal matching between the receptive field responses computed under different viewing conditions for the same scene or for a structurally similar spatio-temporal event.
We develop this theory for the idealized generalized Gaussian derivative model of visual receptive fields, which has been previously demonstrated to rather well model the qualitative shape of simple cells for higher mammals. Formal transformation properties of the receptive fields are stated for the 4 main fundamental classes of primitive geometric image transformations, and it is shown that a visual system, based on such computational primitives, will have the ability to match the spatio-temporal receptive responses computed from dynamic scenes under the variabilities caused by composed variations in the viewing conditions.
We conclude the treatment by discussing and providing potential support for a working hypothesis that the receptive fields of simple cells in the primary visual cortex ought have the shapes of their receptive fields expanded over the degrees of freedom of the corresponding geometric image transformations. We also outline directions for further research, to seek additional characterization and neurophysiological or psychophysical support for the stated predictions, regarding variabilities in the shapes of the spatial or spatio-temporal receptive fields of simple cells in the primary visual cortex.
journal: arXiv preprint
1 Introduction
When to understand the function of the brain, one may ask if there could be overall principles that determine how the neural circuits have evolved to process the information that reach the sensory organs. Specifically, regarding the visual and auditory senses, which provide ways for an observer to infer cues about distant objects in the environment without touching the objects physically, one may ask if there could be special structures in the incoming flows of light and sound waves, that have shaped how the visual and the auditory senses have evolved over the ages of evolution.
Concerning vision, where one of the tasks is to recognize objects from the light that is reflected from them, it is clear that the identity of an object remains, even though we may view the object at different positions in the world, from different distances, different viewing directions and different relative motions between the object and the observer. Concerning the related task of recognizing spatio-temporal events, we may also regard different instantiations of spatio-temporal events, that occur either faster or slower relative to each other, as similar types of spatio-temporal events, under variations of the temporal scales in the time-varying phenomena. Concerning hearing, in a corresponding manner, it is clear that auditory stimuli may carry qualitatively similar information at different time moments, differing in phase and absolute pitch, or differing in the time frame over which the stimuli is presented. Given this basic property, that objects in the environment preserve their identity over classes of geometric transformations, one may hence ask if this property could be regarded as a determining principle for the evolution of visual perception.
The overall theme of this paper is to address this general problem, regarding vision, from the viewpoint of the interactions between geometric image transformations and receptive field responses, specifically manifested in terms of a certain geometric structure on the families of receptive fields corresponding to the simple cells in the primary visual cortex. Formally, we will express these relations in terms of provable covariance [^1] Structurally, the notion of covariance is related to the notion of invariance. The notion of covariance is, however, a weaker and more general property, which means that covariant visual processes may carry more complementary information. Invariant visual processes can, in turn, be constructed on top of covariant visual processes, by adding complementary selection mechanisms over the parameters of covariant processes, for example, the shape parameters of the here considered families of receptive fields. properties of the receptive fields under a specific set of geometric image transformations, to be described further below, based on the following normative arguments, which for a general purpose vision system under sufficient evolutionary pressure could determine how the neural representations would be adapted to structural properties of the stimuli that originate from the environment:
When a visual observer views objects in the environment, the resulting image data on the retina or the image plane in the camera can exhibit a substantial variability, as caused by the geometric image transformations induced by variabilities in the viewing conditions. Specifically, by the variabilities caused by varying (i) the distance, (ii) the viewing direction and (iii) the relative motion between the object and the observer, this will result in geometric image transformations, that to first order of approximation can be modelled in terms of (i) spatial scaling transformations, (ii) spatial affine deformations and (iii) Galilean transformations, see Figure 1 for illustrations. By additionally (iv) viewing different instances of a similar spatio-temporal event that occurs either faster or slower relative to a previously observed reference view, a visual observer can also experience variabilities in terms of (iv) temporal scaling transformations.
Uniform spatial scaling transformations caused by varying the distance between the object and the observer
Refer to caption Refer to caption Refer to caption
Non-isotropic spatial affine transformations caused by varying the viewing direction relative to the object
Refer to caption Refer to caption Refer to caption
Galilean transformations caused by relative motions between objects in the environment and the viewing direction
Refer to caption Refer to caption Refer to caption
Figure 1: Illustrations of variabilities in spatial and spatio-temporal image structures as caused by natural geometric image transformations. (top row) When the distance between the object and the observer is varied, this will lead to perspective image transformations, that to first order of approximation can be modelled as local uniform spatial scaling transformations. (middle row) When the viewing direction is varied relative to the object, this will lead to projective transformations between the two views, that to first order of approximation can be modelled as local spatial affine transformations. (bottom row) When there is relative motion between the object and the observer, the corresponding spatio-temporal image transformations can to first order of approximation be modelled as local Galilean transformations. (Figures in the bottom row reproduced from Lindeberg (2023b) with permission (Open Access).)
When the visual system operates on the resulting spatial or spatio-temporal image data by either purely spatial or joint spatio-temporal receptive fields in the earliest layers of the visual hierarchy, the effects of these types of geometric image transformations will have a very strong influence on the receptive field responses. Despite such huge variabilities occurring regularly when observing natural scenes under generic viewing conditions, we as visual observers are nevertheless able to maintain the identity of objects and to function robustly under such variabilities in the image data.
Given these observations, one may ask if the biological vision systems for different species have evolved to handle the influence of geometric image transformations on the image data, to be able to maintain an identity of the responses from the spatial and spatio-temporal receptive fields under the huge variabilities in image data that can be generated from similar scenes as depending on variabilities in the viewing conditions.
The subject of this paper is to give both (i) an overview, (ii) a set of conceptual extensions, and (iii) a set of biological implications and predictions from a theory that has been developed to address this topic in a recent series of papers in Lindeberg (2021b, 2023b, 2025d, 2025a, 2025c, 2025b, 2025e), and which constitutes a substantially extended version of an earlier prototype to this theory in Lindeberg (2011, 2013a). Compared to the original papers, the presentation in this review will be simplified, with less focus on the mathematical details and more emphasis on the main ideas and concepts, thus aimed at making overall results from the proposed theory more easily accessible for a wider audience.
Compared to the previous publications on this topic, this presentation will be significantly more integrated, where the different types of variabilities that were addressed individually in in Lindeberg (2025c; 2025a) are here addressed in a unified manner, thus reframing the individually addressed variabilities in Lindeberg (2025c; 2025a) into a wider context. In relation to the early approach to such variabilities in Lindeberg (2023b), the resulting biological hypotheses will also be formulated based on significantly further theoretical depth, [^2] Beyond the first contributions in Lindeberg (2025c; 2025a), the theoretical background to the biological hypotheses in this paper will be based on the set of affine covariant derivative-based models for visual receptive fields to be described in Section 5.3.2, that were published after the publications of the earlier more biologically oriented contributions, and which will be built upon in the formulation of the more explicit affine-covariant models of visual receptive fields in Section 5.4 and 5.5, will then constitute a more solid foundation for the extended biological predictions to be presented in Section 6. In addition, we provide an alternative model for handling variabilities with respect to uniform spatial scaling transformations and temporal scaling transformations in Section 6, by only representing the finest spatial and temporal scales in the sensorium, as enabled by a formal theory for cascade smoothing properties, and as developed in the recent companion paper (Lindeberg 2025g). as enabled by the generalized mathematical theory presented in Lindeberg (2025d; 2025g).
We will additionally describe a set of significant extensions [^3] The explicit statements in Sections 5.4, 5.5 and 6 go significantly further compared to the related results in the previous publications. relative to the previously presented theoretical results, as enabled by describing the overall theory in a unified manner relative to the previous more specialized technical contributions. Specifically, we will use results from this theory to address the topic of variabilities of simple cells [^4] According to the taxonomy of neurons in the primary visual cortex by Hubel and Wiesel (1959; 1962; 1968; 2005), a visual neuron is said to be simple if: (i) it has distinct excitatory and inhibitory subregions, (ii) it obeys roughly linear summation properties and (iii) the excitatory and inhibitory regions balance each other in diffuse lighting. in the primary visual cortex, based on a hypothesis that the shapes of the spatial or spatio-temporal receptive fields ought to span the degrees of the geometric image transformations that are involved in the image formation process, as well as provide further depth to the alternative working hypothesis that a vision system could choose to only implement the lowest levels of spatial and temporal scales in the sensorium corresponding to the simple cells. This will be done by making use of cascade smoothing properties over spatial and temporal scales, as enabled by the recently developed hybrid Lie semi-group theory over the filter shape parameters in the multi-parameter generalized Gaussian derivative model for visual receptive fields, presented in Lindeberg (2025g).
Based on these theoretical results, we will then propose an extended set of theoretical predictions in Section 6, and relate them to the neurophysiological results that we currently find as available, and also propose new neurophysiological and psychophysical experiments to address the questions posed from this theoretical study, and which previous neurophysiological and psychophysical experiments may not have been designed to address.
A main fundament of the overall presented theoretical framework is to postulate an identity between the receptive field responses computed from different observations of the same scene or the same spatio-temporal event, by requiring the family of receptive fields to be covariant under local linearizations of the considered classes of geometric image transformations. Covariance, also referred to as equivariance in some literature, essentially means that the families of receptive fields are to be well-behaved under the corresponding classes of geometric image transformations, in such a way that the result of computing a receptive field response from geometrically transformed image data should correspond to applying the same type of geometric transformation to the receptive field response of the original image data before the image transformation.
In this way, the notion of covariance makes it possible to establish a notion of identity between the receptive field responses computed from sets of image data that have been transformed by the geometric image transformations, as induced by varying the viewing conditions in terms of the distance, the viewing direction, the relative motion between objects in the scene and observer, and also the image transformations induced by viewing similar types of spatio-temporal events that may occur either faster or slower relative to a previously observed reference view.
Specifically, the notion of covariance makes it possible to perfectly match the receptive field responses between different views of the same scene or spatio-temporal event, in such a way that the receptive field responses are either exactly equal or very similar for local linearizations of the geometric image transformations, for a particular way of formulating the receptive fields corresponding to idealized models of simple cells according to the considered generalized Gaussian derivative model for visual receptive fields.
A more general underlying biological motivation to this study is that, while it is well-known that neurons in the brain tend to exhibit different types of variabilities, it may on the other hand be usually less known what are the sources to those variabilities. See also Młynarski (2025) for a recent perspective of this, regarding the retina and of explaining the functional variability in the retina over the dimensions of sensory coding, stimulus and noise statistics and collective coding over neuronal populations, with special emphasis on the need for normative theories. In this treatment, we present a systematic and theoretically principled study, that predicts a set of variabilities regarding the shapes of the receptive fields of simple cells in the primary cortex, based on a normative theory of visual receptive fields, and with a theoretically well-founded explanation in terms of covariance properties under geometric image transformations of the receptive fields in the lower layers of the visual hierarchy, to in turn enable the computation of invariant visual representations with regard to the influence of geometric image transformations at higher layers in the visual pathway.
1.1 Structure of this presentation
In this paper, we will summarize the main components of this theory, infer biological interpretations as well as state biological predictions of the theory in the following way: After a brief overview of related work in Section 2, Section 3 starts by first defining the main classes of geometric image transformations that we consider.
Then, Section 4 defines the axiomatically determined idealized models for visual receptive fields that we base this work upon, complemented in Section 5 by describing the covariance properties of the idealized receptive fields under the considered classes of geometric image transformations. Section 6 then addresses whether we can regard the spatial and the spatio-temporal shapes of simple cells in the primary visual cortex of higher mammals to span the variabilities of geometric image transformations, to support explicitly covariant families of visual receptive fields. Section 6 also describes ideas to future neurophysiological and psychophysical experiments to investigate this topic in more detail. Finally, Section 7 concludes with a summary and discussion.
As a guide to the reader, this presentation is aimed at both researchers interested in theoretical and computational modelling of visual receptive fields and researchers interested in characterizing the properties of the visual neurons in the visual pathway, including neurophysiological and psychophysical experimentalists. For readers without a strong mathematical background, a few of the sections in the main Section 5 in this paper may by necessity be somewhat technical, to make it possible to reproduce the main ideas, concepts and implications from the underlying mathematical theory. For a reader more interested in the biological interpretations and implications, a shorter path should be possible, by first reading the introductory Sections 3 and 4 concerning the image geometry and the receptive field models and then proceeding directly to the treatment in Section 6, about whether the shapes of the receptive fields of simple cells in the primary visual cortex can be regarded as spanning the degrees of freedom of the set of primitive geometric image transformations. For filling in possible missing complementary details, a good start could then be to backtrack from the references to specific sections and equations from the summary and discussion in Section 7.
2 Relations to previous work
Concerning variabilities of image data under spatial scaling transformations, there are several sources of evidence that demonstrate scale-invariant processing in the primate visual cortex; see Biederman and Cooper (1992), Logothetis et al. (1995), Ito et al. (1995), Furmanski and Engel (2000), Hung et al. (2005) and Isik et al. (2013).
Given that scale-covariant image operations in the lower layers constitute a powerful precursor to scale-invariant image operations in higher layers in the visual hierarchy (see Lindeberg (2021b) Appendix I), one may hence ask if the earliest layers of the visual system of higher mammals can be regarded as able to process the image data in a way that can be modelled in terms of scale covariance. In a corresponding manner, one may ask if such a covariance property would also extend to other types of geometric image transformations, in relation to viewing objects, scenes and spatio-temporal events under different types of viewing conditions. Notably, the correspondence of invariance principles has been postulated for biological vision systems by Rolls (1994), DiCarlo and Maunsell (2000), Grimes and Rao (2005), Quiroga et al. (2005), DiCarlo and Cox (2007), and Goris and Beek (2009) and Lindeberg (2013b).
Our knowledge about the functional properties of the receptive fields of simple cells in the primary visual cortex originates from the pioneering work by Hubel and Wiesel (1959; 1962; 1968; 2005) followed by more detailed characterizations by DeAngelis et al. (1995; 2004), Ringach (2002; 2004), Conway and Livingstone (2006), Johnson et al. (2008), Walker et al. (2019) and De and Horwitz (2021).
Computational models of simple cells have specifically been expressed in terms of Gabor filters by Marcelja (1980), Jones and Palmer (1987a; 1987b), Porat and Zeevi (1988), Ringach (2002; 2004), Serre et al. (2007) and De and Horwitz (2021), and in terms of Gaussian derivatives by Koenderink and van Doorn (1984; 1987; 1992), Young (1987), Young et al. (2001; 2001) and Lindeberg (2013a; 2021b). Theoretical models of early visual processes have also been formulated based on Gaussian derivatives by Lowe (2000), May and Georgeson (2007), Hesse and Georgeson (2005), Georgeson et al. (2007), Hansen and Neumann (2008), Wallis and Georgeson (2009), Wang and Spratling (2016), Pei et al. (2016), Ghodrati et al. (2017), Kristensen and Sandberg (2021), Abballe and Asari (2022), Ruslim et al. (2023) and Wendt and Faul (2024).
Learning-based schemes to model visual receptive fields from training data have also been proposed by Rao and Ballard (1998), Olshausen and Field (1996; 1997), Simoncelli and Olshausen (2001), Geisler (2008), Hyvärinen et al. (2009), Lörincz et al. (2012) and Singer et al. (2018). Poggio and Anselmi (2016) did on the other hand propose to model learning of invariant receptive fields based on group theory. More recently, deep learning approaches have been applied for modelling visual receptive fields (Keshishian et al. 2020), although one may also raise issues concerning the applicability of such approaches, see Bae et al. (2021), Bowers et al. (2022), Heinke et al. (2022), Wichmann and Geirhos (2023) and the references therein.
The main subject of this paper is to model the receptive fields of simple cells based on the normative theory for visual receptive fields proposed in Lindeberg (2021b) in terms of the generalized Gaussian derivative model, and then consider the influence on the resulting receptive field responses caused by variabilities in geometric image transformations.
This approach does specifically have structural similarities to the recently developed area of geometric deep learning (Bronstein et al. 2021, Gerken et al. 2023), where deep networks are formulated from the constraint that they should be well-behaved under the influence of geometric image transformations. For examples of deep networks that are covariant under spatial scaling transformations, see Worrall and Welling (2019), Bekkers (2020), Sosnovik et al. (2020, 2021), Zhu et al. (2022), Jansson and Lindeberg (2022), Lindeberg (2022), Zhan et al. (2022), Wimmer et al. (2023) and Perzanowski and Lindeberg (2025).
3 Main classes of locally linearized geometric image transformations
For a monocular observer, that views the objects in a 3-D scene by a planar 2-D image sensor, the projection is described by a non-linear perspective projection model. For a binocular observer or multiple monocular observers, that view the same 3-D scene from multiple observation points and multiple viewing directions, the transformations between the different views of the same scene are described by non-linear projective transformations. To substantially simplify these non-linear projection models, we will linearize them locally around each point in terms of local first-order derivatives, which will then result in the following classes of linear projection models applied to the image coordinates of the form $x=(x_{1},x_{2})^{T}\in{\mathbb{R}}^{2}$ and the temporal variable $t\in{\mathbb{R}}$:
Uniform spatial scaling transformations:
$f^{\prime}(x^{\prime})=f(x)\quad\quad\mbox{for}\quad\quad x^{\prime}=S_{x}\,x,$ (1)
where $S_{x}\in{\mathbb{R}}_{+}$ is a spatial scaling factor.
Spatial affine transformations:
$f^{\prime}(x^{\prime})=f(x)\quad\quad\mbox{for}\quad\quad x^{\prime}=A\,x,$ (2)
where $A$ is a non-singular $2\times 2$ matrix with strictly positive eigenvalues.
Galilean transformations:
$f^{\prime}(x^{\prime},t^{\prime})=f(x,t)\quad\quad\mbox{for}\quad\quad x^{\prime}=x+u\,t,\;t^{\prime}=t,$ (3)
where $u=(u_{1},u_{2})^{T}\in{\mathbb{R}}^{2}$ is a 2-D velocity vector.
Temporal scaling transformations:
$f^{\prime}(x^{\prime},t^{\prime})=f(x,t)\quad\quad\mbox{for}\quad\quad t^{\prime}=S_{t}\,t,\;x^{\prime}=x,$ (4)
where $S_{t}\in{\mathbb{R}}_{+}$ is a temporal scaling factor.
Of particular interest is to compose these geometric transformations in the following way when observing dynamic scenes with either monocular or binocular locally linearized camera models (Lindeberg 2025d Equations (222)–(223)):
$\begin{split}x^{\prime}=S_{x}\,(A\,x+u\,t),\end{split}$ (5)
$\begin{split}t^{\prime}=S_{t}\,t.\end{split}$ (6)
Figure 2: Illustration of the geometry underlying the composed locally linearized projection model in Equations (5) and (6) for a single monocular view. Here, a local, possibly moving, surface patch is projected to an arbitrary view indexed by $k$ in a multi-view locally linearized projection model, with the fixation point $F$ on the surface mapped to the origin $O^{(k)}=0$ in the image plane for the observer with the optic center $P^{(k)}$. Then, any point in the tangent plane to the surface at the fixation point, as parameterized by the local coordinates $\xi$ in a coordinate frame attached to the tangent plane of the surface with $\xi=0$ at the fixation point $F$, is by the local linearization mapped to the image point $x^{(k)}$. (Figure reproduced from Lindeberg (2025d) with permission (OpenAccess).)
Figure 3: Illustration of the underlying geometric situtation for the locally linearized transformations between pairwise views of the same, possibly moving, local surface patch according to Equation (9). Here, the view indexed by $\tilde{k}$ constitutes the reference view and the view indexed by $k$ constitutes an arbitrary view. By a connection of the point $\xi$ being the same in two instances of Figure 3, we can from the parameters $S_{t}$, $A$ and $u$ of the monocular mappings for optical centers based on the indices $k$ and $\tilde{k}$ establish a relationship between the matching image points $x^{(k)}$ and $x^{(\tilde{k})}$ in these two views. (Figure reproduced from Lindeberg (2025d) with permission (OpenAccess).)
Then, specifically
  • the $2\times 2$ affine transformation matrix $A$ models the orthonormal projection of surface patterns from the tangent plane of a local surface patch to a plane, parallel with the image plane of the observer,
  • the velocity vector $u=(u_{1},u_{2})^{T}\in{\mathbb{R}}^{2}$ models the projection of the 3-D motion vector $U=(U_{1},U_{2},U_{3})^{T}$ of local surface patterns onto a plane, parallel to the image plane, by local orthonormal projection,
  • the spatial scaling factor $S_{x}\in{\mathbb{R}}_{+}$ models the perspective scaling factor proportional to the inverse depth $Z$, which will then affect both the projection of a spatial surface pattern and the magnitude of the perceived motion in the image plane, and
  • the temporal scaling factor $S_{t}\in{\mathbb{R}}_{+}$ models the variability of similar spatio-temporal events, that may occur either faster or slower, when observing different instances of a similar event at different occasions.
Thereby, the composed image transformation model captures the variabilities of the scaled orthographic projection model, complemented with a variability over projections of 3-D motions between an observed object and the observer, including spatio-temporal events, that may occur faster or slower relative to a reference view, see Figure 3 for an illustration.
By further considering a pair of such projection equations for the indices $k$ and $\tilde{k}$ of the observation points, and introducing the alternative parameterizations of the parameters according to Lindeberg (2025d) Equations (308)–(309)
$\tilde{B}^{(k)}=\frac{S_{x}^{(k)}}{S_{x}^{(\tilde{k})}}\,A^{(k)}(A^{(\tilde{k})})^{-1}$ (7)
and
[…]
we have that corresponding image points $x^{(k)}$ and $x^{(\tilde{k})}$ between these views can be expressed as (Lindeberg 2025d Equation (299))
$x^{(k)}=\tilde{B}^{(k)}\,x^{(\tilde{k})}+\tilde{u}^{(k)}\,t,$ (9)
where
  • $x^{(k)}\in{\mathbb{R}}^{2}$ is the locally linearized projection of the physical point on the surface pattern in the view from the observer with index $k$ at time $t$,
  • $x^{(\tilde{k})}\in{\mathbb{R}}^{2}$ is the locally linearized projection of the physical point on the surface pattern in the view from the observer with index $\tilde{k}$ at time $t$,
  • $\tilde{B}^{(k)}$ is a non-singular $2\times 2$ affine projection matrix for the observer with index $k$ in relation to an observation from a reference view with index $\tilde{k}$, and
  • $\tilde{u}^{(k)}\in{\mathbb{R}}^{2}$ is a corresponding 2-D relative motion vector for the observer with index $k$ in relation to an observation from a reference view with index $\tilde{k}$,
see Figure 3 for an illustration.
In these ways, we can based on the four primitive geometric image transformations according to Equations (1)–(4) model both locally linearized monocular perspective projections and locally linearized binocular projective projections of dynamic scenes, based on joint compositions of these primitives according to Equations (5), (6) and (9).
$\sigma_{x}=1$ $\sigma_{x}=\sqrt{2}$ $\sigma_{x}=2$ $\sigma_{x}=2\sqrt{2}$ $\sigma_{x}=4$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Illustration of the variability of spatial receptive fields under uniform spatial scaling transformations. Here, the first-order directional derivative of the Gaussian kernel $T_{\varphi}(x;\;s,\Sigma)=\partial_{\varphi}(g(x;\;s,\Sigma))$ in the horizontal direction $\varphi=0$ is shown for different values of the spatial scale parameter $\sigma_{x}=\sqrt{s}$ for the special case of using an isotropic spatial covariant matrix with $\Sigma=I$. The variability of this spatial scale parameter makes it possible to handle objects of different size in the world as well as objects at different distances to the camera. (Horizontal axes: spatial coordinate $x_{1}\in[-10,10]$. Vertical axes: spatial coordinate $x_{2}\in[-10,10]$.)
Figure 5: Illustration of the variability of spatial receptive fields under non-isotropic spatial affine transformations. Here, first-order directional spatial derivatives of Gaussian kernels $T_{\varphi}(x;\;s,\Sigma)=\partial_{\varphi}(g(x;\;s,\Sigma))$ are shown under variations of the preferred image orientation and the degree of elongation of receptive fields of the spatial covariance matrices $\Sigma$, corresponding to a uniform distribution on a hemisphere. The variability of the spatial covariance matrix $\Sigma$ corresponds to varying the slant and the tilt angles of the surface patch over all the angles on the visible hemisphere. (In this figure, the spatial scale parameters of the receptive fields have been normalized, such that the maximum eigenvalue of the spatial covariance matrix $\Sigma$ is the same for all the receptive fields.) (Horizontal and vertical axes: the spatial coordinates $x_{1}$ and $x_{2}$, for multiple spatial receptive fields shown within the same frame.)
v=1v=-1 v=1/2v=-1/2 $v=0$ $v=1/2$ $v=1$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: Illustration of the variability of spatio-temporal receptive fields under Galilean transformations. Here, the mixed spatio-temporal derivative $T_{x{\bar{t}}}(x,t;\;s,\tau,v)=\partial_{x}\,\partial_{\bar{t}}\left(g(x-v\,t;\;s)\,\Psi(t;\;\tau,c)\right)$ over a 1+1-D spatio-temporal domain is shown for different values of the velocity parameter $v$, based on using a first-order Gaussian derivative over the spatial domain and a first-order derivative of the time-causal limit kernel over the temporal domain, for $s=\sigma_{x}^{2}$, $\tau=\sigma_{t}^{2}$ and $c=2$ with $\sigma_{x}=1/2$ and $\sigma_{t}=1$. This variability corresponds to varying the relative motion between the viewing direction and a moving local surface patch. (Horizontal axes: spatial coordinate $x\in[-3.5,3.5]$. Vertical axes: time $t\in[0,3.5]$.)
$\sigma_{t}=1$ $\sigma_{t}=2$ $\sigma_{t}=4$
Refer to caption Refer to caption Refer to caption
Figure 7: Illustration of the variability of purely temporal receptive fields under temporal scaling transformations. Here, the first-order scale-normalized temporal derivative of the time-causal limit kernel $T_{t}(t;\;\tau)=\sqrt{\tau}\,\partial_{t}\left(\Psi(t;\;\tau,c)\right)$ for $c=2$ is shown for different values of the temporal scale parameter $\sigma_{t}=\sqrt{\tau}$. This variability corresponds to observing temporal structures that occur either faster or slower relative to a previously observed reference view. (Horizontal axes: time $t\in[0,16]$. Vertical axes: magnitude of the scale-normalized derivative $\in[-0.5,1]$.)
4 Idealized receptive fields according to the generalized Gaussian derivative model for visual receptive fields
In this section, we will review main components in the generalized Gaussian derivative theory for visual receptive fields, as arising from the normative theory for visual receptive fields in Lindeberg (2021b).
4.1 Receptive field models in terms of linear spatial or spatio-temporal convolution operations
Given spatial image data $f\colon{\mathbb{R}}^{2}\rightarrow{\mathbb{R}}$ expressed on the form $f(x)=f(x_{1},x_{2})$ for the image coordinates $x=(x_{1},x_{2})^{T}\in{\mathbb{R}}^{2}$ or spatio-temporal image data $f\colon{\mathbb{R}}^{2}\times{\mathbb{R}}\rightarrow{\mathbb{R}}$ expressed on the form $f(x,t)=f(x_{1},x_{2},t)$ with an additional dependency on the temporal variable $t\in{\mathbb{R}}$, a (linear) spatial receptive field $T\colon{\mathbb{R}}^{2}\rightarrow{\mathbb{R}}$ or a (linear) spatio-temporal receptive field $T\colon{\mathbb{R}}^{2}\times{\mathbb{R}}\rightarrow{\mathbb{R}}$ can be seen as a spatial or a spatio-temporal convolution kernel, that is to be applied to the image data $f$ according to
$(T*f)(x)=\int_{\xi\in{\mathbb{R}}^{2}}T(\xi)\,f(x-\xi)\,d\xi$ (10)
in the case of a purely spatial image domain, or according to
$(T*f)(x,t)=\int_{\xi\in{\mathbb{R}}^{2}}\int_{\eta\in{\mathbb{R}}}T(\xi,\eta)\,f(x-\xi,t-\eta)\,d\xi\,d\eta$ (11)
in the case of a joint spatio-temporal image domain.
4.2 Covariance properties of spatial and spatio-temporal receptive responses under spatial translations in the image plane and temporal shifts
Because of this convolution structure, the receptive field responses are covariant under spatial translations in the image plane according to
$f^{\prime}(x^{\prime})=f(x)\quad\quad\mbox{or}\quad\quad f^{\prime}(x^{\prime},t^{\prime})=f(x,t)$ (12)
for
$x^{\prime}=x+\Delta x\quad\quad\mbox{where}\quad\quad\Delta x\in{\mathbb{R}}^{2}$ (13)
and $t^{\prime}=t$, in the sense that the corresponding spatial or spatio-temporal receptive field responses L=TfL=T*f and $L^{\prime}=T*f^{\prime}$ then satisfy
$L^{\prime}(x^{\prime})=L(x)\quad\quad\mbox{or}\quad\quad L^{\prime}(x^{\prime},t^{\prime})=L(x,t).$ (14)
Similarly, under a temporal shift of spatio-temporal image data of the form
$f^{\prime}(x^{\prime},t^{\prime})=f(x,t)$ (15)
for
$t^{\prime}=t+\Delta t\quad\quad\mbox{where}\quad\quad\Delta t\in{\mathbb{R}}$ (16)
and $x^{\prime}=x$, the corresponding spatio-temporal receptive field responses L=TfL=T*f and $L^{\prime}=T*f^{\prime}$ satisfy
$L^{\prime}(x^{\prime},t^{\prime})=L(x,t).$ (17)
Because of these covariance properties, a vision system based on receptive field responses, that can be modelled in terms of convolution operations, will handle objects at different positions [^5] In this treatment, we disregard the effects of a spatially varying sampling density of the receptive fields on a foveated sensor, such as the primate retina. For a principled treatment of such spatial sampling effects with respect to the receptive field responses, see Lindeberg and Florack (1994), with a condensed summary of some of the main results in Lindeberg (2013a) Section 7. in the image plane as well as temporal events, that occur at different time moments, in a similar manner.
A main subject of this paper is to present a set of theoretical extensions to this linear convolution structure, to make it possible for receptive fields with structurally similar properties as the simple cells in the primary visual cortex to handle more developed sets of geometric image transformations applied to the image data used as input to a vision system.
4.3 Idealized spatial or spatio-temporal models for simple cells in the primary visual cortex
To handle the additional influence on the receptive fields, due to the in Section 3 described classes of geometric image transformations, we will consider idealized models for simple cells, based on the generalized Gaussian derivative model for visual receptive fields, as initiated in the early work in Lindeberg (2011; 2013a) and then further refined regarding the temporal domain in Lindeberg (2016; 2021b).
The receptive fields according to this model have been obtained based on axiomatic derivations that reflect symmetry properties of the environment in combination with internal consistency requirements to guarantee theoretically well-founded treatment of image structures over different spatial and temporal scales. In this respect, the families of receptive fields have been formulated in a theoretically well-founded manner.
According to the underlying normative theory for visual receptive fields, the shapes of the receptive fields are parameterized by a set of filter parameters, with linear models of purely spatial receptive fields corresponding to simple cells formulated in terms of affine Gaussian derivatives of the form
$T_{\text{simple}}(x_{1},x_{2};\;\sigma_{\varphi},\varphi,\Sigma_{\varphi},m)=\\ =T_{\varphi^{m},\scriptsize\mbox{norm}}(x_{1},x_{2};\;\sigma_{\varphi},\Sigma_{\varphi})=\sigma_{\varphi}^{m}\,\partial_{\varphi}^{m}\left(g(x_{1},x_{2};\;\Sigma_{\varphi})\right),$ (18)
where
  • $\varphi\in[-\pi,\pi]$ is the preferred orientation of the receptive field,
  • $\sigma_{\varphi}\in{\mathbb{R}}_{+}$ is the amount of spatial smoothing (in units of the spatial standard deviation),
  • $\partial_{\varphi}^{m}=(\cos\varphi\,\partial_{x_{1}}+\sin\varphi\,\partial_{x_{2}})^{m}$ is an $m$:th-order directional derivative operator in the direction $\varphi$,
  • $\Sigma_{\varphi}$ is a $2\times 2$ symmetric positive definite covariance matrix, with one of its eigenvectors in the direction of $\varphi$,
  • $g(x;\;\Sigma_{\varphi})$ is a 2-D affine Gaussian kernel, with its shape determined by the spatial covariance matrix $\Sigma_{\varphi}$
    $g(x;\;\Sigma_{\varphi})=\frac{1}{2\pi\sqrt{\det\Sigma_{\varphi}}}e^{-x^{T}\Sigma_{\varphi}^{-1}x/2}$ (19)
    for $x=(x_{1},x_{2})^{T}\in{\mathbb{R}}^{2}$.
Concerning time-dependent image data, spatio-temporal receptive fields corresponding to simple cells are, in turn, formulated according to
$\begin{split}T_{\text{simple}}(x_{1},x_{2},t;\;\sigma_{\varphi},\sigma_{t},\varphi,v,\Sigma_{\varphi},m,n)\end{split}$
$\begin{split}&=T_{{\varphi}^{m},{\bar{t}}^{n},\scriptsize\mbox{norm}}(x_{1},x_{2},t;\;\sigma_{\varphi},\sigma_{t},v,\Sigma_{\varphi})\end{split}$
$\begin{split}&=\sigma_{\varphi}^{m}\,\sigma_{t}^{n}\,\partial_{\varphi}^{m}\,\partial_{\bar{t}}^{n}\left(g(x_{1}-v_{1}t,x_{2}-v_{2}t;\;\Sigma_{\varphi})\,h(t;\;\sigma_{t})\right),\end{split}$ (20)
where for the symbols not previously defined in connection with Equation (18) we have that:
  • $\sigma_{t}$ represents the amount of temporal smoothing (in units of the temporal standard deviation),
  • $v=(v_{1},v_{2})^{T}$ represents a local motion vector in the direction $\varphi$ of the spatial orientation of the receptive field,
  • $\partial_{\bar{t}}^{n}=(\partial_{t}+v_{1}\,\partial_{x_{1}}+v_{2}\,\partial_{x_{2}})^{n}$ represents an $n$:th-order velocity-adapted temporal derivative operator, and
  • $h(t;\;\sigma_{t})$ represents a temporal smoothing kernel with temporal standard deviation $\sigma_{t}$.
For the case of the temporal domain being non-causal (meaning that the future relative to an temporal moment can be accessed, as it can be on pre-recorded video data), the temporal kernel can be chosen as the 1-D Gaussian kernel
$h(t;\;\sigma_{t})=\frac{1}{\sqrt{2\pi}\sigma_{t}}\,e^{-t^{2}/2\sigma_{t}^{2}},$ (21)
whereas in the case of the temporal domain being time-causal (implying the more realistic real-time scenario, where the future cannot be accessed), the temporal kernel can determined as the time-causal limit kernel (Lindeberg 2016 Section 5; Lindeberg 2023a Section 3)
$h(t;\;\sigma_{t})=\psi(t;\;\sigma_{t},c),$ (22)
characterized by having a Fourier transform of the form
$\hat{\Psi}(\omega;\;\sigma_{t},c)=\prod_{k=1}^{\infty}\frac{1}{1+i\,c^{-k}\sqrt{c^{2}-1}\,\sigma_{t}\,\omega}.$ (23)
This form of the temporal smoothing function corresponds to using an infinite set of first-order integrators that are coupled in cascade, with the time constants chosen so as to specifically enable temporal scale covariance. The distribution parameter $c>1$, often chosen as $c=\sqrt{2}$ or $c=2$, in this temporal smoothing function reflects the ratio between adjacent discrete temporal scale levels in the corresponding temporal scale-space model.
In Lindeberg (2021b), it was demonstrated that idealized receptive field models of these types do rather well model the qualitative shape of biological simple cells as obtained by neurophysiological measurements by DeAngelis et al. (1995; 2004), Conway and Livingstone (2006) and Johnson et al. (2008); see Figures 12–18 in Lindeberg (2021b) for comparisons between biological receptive fields and idealized models thereof, based on the generalized Gaussian derivative model for visual receptive fields.
Specifically, this formulation of the idealized models of spatial and spatio-temporal receptive fields implies that the shapes of the receptive fields are expanded with respect to the degrees of freedom of the corresponding geometric image transformations, as illustrated in Figures 47.
(a) Non-covariant receptive fields (b) Covariant receptive fields
Refer to caption Refer to caption
Figure 8: Illustration of the importance of covariance properties of the receptive fields when computing receptive field responses of a scene under different viewing conditions. (left) If the receptive field family is not covariant with respect to the appropriate class of geometric image transformations, then the backprojections of the receptive fields onto the tangent plane of an observed local surface patch will, in general, be different in the tangent plane of the surface. Thereby, if the receptive field responses are to be used for, to example, for computing local shape properties of the surface patch, then those shape estimates may be strongly biased because of effects of that mismatch between the backprojected receptive fields. (right) If the receptive field family is covariant with respect to the relevant class of geometric image transformations, then it is, on the other hand, possible to match the parameters of the receptive fields over the two image domains in such a way that the backprojected receptive fields do to first order of approximation coincide in the tangent plane of the surface. Thereby, the source to bias caused by a mismatch of the backprojected receptive fields can be substantially reduced, which enables to computation of more accurate estimates of the local surface shape. While this example concerns spatial receptive fields corresponding to two views with different viewing directions relative to a static scene, corresponding effects regarding the backprojections of the receptive fields will occur also when computing spatio-temporal receptive field responses for dynamic scenes. (Figure reproduced from Lindeberg 2023b with permission (Open Access).)
5 Covariance properties of idealized models of simple cells under locally linearized geometric image transformations
The notion of covariance, in some literature also referred to as equivariance, means that the family of receptive fields is to be well-behaved under a given class of geometric image transformations $\cal G$}, see Figure 8 for an illustration.
[…]
${\cal R}\,{\cal G}\,f={\cal G}\,\tilde{{\cal R}}\,f.$ (24)
[…]
A very useful property of the receptive fields according the generalized Gaussian derivative model for visual receptive fields is that (see Lindeberg (2023b) for details):
  • for both the purely spatial model (18) for simple cells and the joint spatio-temporal model (4.3) for simple cells, the receptive field responses are covariant under both uniform spatial scaling transformations of the form (1) and spatial affine transformations of the form (2), and
  • the joint spatio-temporal model (4.3) for simple cells is also covariant under Galilean transformations of the form (3) and temporal scaling transformations of the form (4).
In the case of using the non-causal Gaussian kernel (21) as the temporal smoothing kernel in the idealized receptive field family, the temporal scale covariance property holds for all non-negative temporal scaling factors $S_{t}\in{\mathbb{R}}_{+}$. In the case of using the time-causal limit kernel (22) as the temporal smoothing kernel, for which the temporal scaling factors do not form a continuum but are discrete, the temporal covariance property holds for all temporal scaling factors $S_{t}$ that are integer powers of the distribution parameter $c$ of the temporal smoothing kernel, according to $S_{t}=c^{i}$ for $i\in{\mathbb{Z}}$.
These properties do thus imply that the receptive fields according to generalized Gaussian derivative model for visual receptive field are well-behaved under both (i) uniform spatial scaling transformations, (ii) spatial affine transformations, (iii) Galilean transformations, and (iv) temporal scaling transformations. In this way, the generalized Gaussian derivative model can consistently process both purely spatial and joint spatio-temporal image data, that are subject to these individual geometric image transformations, as well as to joint combinations thereof.
5.1 Formal statement of the covariance properties for the pure spatial and spatio-temporal smoothing operations without spatial or temporal differentiation
To express the joint covariance property in a compact manner, let us consider the composed geometric transformation of the form in Equations (5)–(6), which models the joint effect of the four types of primitive geometric image transformations (1)–(4) when observing a possibly moving local surface patch with scaled orthographic projection, from possibly different viewing distances and viewing directions, in situations where there could be relative motions between the object and the observer, and also a spatio-temporal event may occur either faster or slower relative to a reference view.
Let us initially disregard the effects of the spatial and the temporal derivative operators in the idealized spatio-temporal receptive field model (18) by setting the differentiation orders to $m=0$ and $n=0$, leading to the following form for the spatio-temporal smoothing kernel [^6] In this treatment, ${\mathbb{S}}_{+}^{2}$ denotes the set of symmetric positive definite $2\times 2$ matrices. $T\colon{\mathbb{R}}^{2}\times{\mathbb{R}}\times{\mathbb{R}}_{+}\times{\mathbb{S}}_{+}^{2}\times{\mathbb{R}}_{+}\times{\mathbb{R}}^{2}\rightarrow{\mathbb{R}}$ according to
$T(x,t;\;s,\Sigma,\tau,v)=g(x-v\,t;\;s,\Sigma)\,h(t;\;\tau)$ (25)
for the alternative parameterization of the spatial and temporal scale parameters according to $s=\sigma_{x}^{2}$ and $\tau=\sigma_{t}^{2}$, where we have here also for forthcoming use redefined the spatial affine Gaussian kernel into the following overparameterized form
$g(x;\;s,\Sigma)=\frac{1}{2\pi\,s\sqrt{\det\Sigma}}\,e^{-x^{T}\Sigma^{-1}x/2s},$ (26)
in order to later more clearly be able to separate the degrees of freedom between pure uniform spatial scaling transformations and more general spatial affine transformations.
Let us also redefine the temporal smoothing kernel $h\colon{\mathbb{R}}\times{\mathbb{R}}_{+}\rightarrow{\mathbb{R}}$ as either the non-causal 1-D Gaussian kernel according to
$h(t;\;\tau)=\frac{1}{\sqrt{2\pi}\sqrt{\tau}}\,e^{-t^{2}/2\tau},$ (27)
or the time-causal limit kernel according to (Lindeberg 2016 Section 5; Lindeberg 2023a Section 3)
$h(t;\;\tau)=\psi(t;\;\tau,c),$ (28)
characterized by having a Fourier transform of the form
$\hat{\Psi}(\omega;\;\tau,c)=\prod_{k=1}^{\infty}\frac{1}{1+i\,c^{-k}\sqrt{c^{2}-1}\,\sqrt{\tau}\,\omega},$ (29)
where we will for all forthcoming use parameterize these kernels in terms of the temporal variance $\tau=\sigma_{t}^{2}$, opposed to instead using the temporal standard deviation $\sigma_{t}$ in Equations (21)–(23).
Next, let us for any 2+1-D spatio-temporal image data $f\colon{\mathbb{R}}^{2}\times{\mathbb{R}}\rightarrow{\mathbb{R}}$ define spatio-temporally smoothed image data $L\colon{\mathbb{R}}^{2}\times{\mathbb{R}}\times{\mathbb{R}}_{+}\times{\mathbb{S}}_{+}^{2}\times{\mathbb{R}}_{+}\times{\mathbb{R}}^{2}\rightarrow{\mathbb{R}}$ according to (Lindeberg 2025d Equation (177))
$L(\cdot,\cdot;\;s,\Sigma,\tau,v)=T(\cdot,\cdot;\;s,\Sigma,\tau,v)*f(\cdot,\cdot).$ (30)
Let us also for geometrically transformed image data
$f^{\prime}(x^{\prime},t^{\prime})=f(x,t)$ (31)
under the composed geometric image transformation according to (5)–(6) define correspondingly spatio-temporally transformed image data according to
$L(\cdot,\cdot;\;s^{\prime},\Sigma^{\prime},\tau^{\prime},v^{\prime})=T(\cdot,\cdot;\;s^{\prime},\Sigma^{\prime},\tau^{\prime},v^{\prime})*f^{\prime}(\cdot,\cdot).$ (32)
Then, as shown in Section 5.2 in Lindeberg (2025d), it holds that these spatio-temporally smoothed image data are equal under the composed geometric image transformation of the form (5)–(6) (Lindeberg 2025d Equation (251))
$L^{\prime}(x^{\prime},t^{\prime};\;s^{\prime},\Sigma^{\prime},\tau^{\prime},v^{\prime})=L(x,t;\;s,\Sigma,\tau,v),$ (33)
provided that the parameters of the receptive fields over the two spatio-temporal image domains are related according to (Lindeberg 2025d Equations (252)–(255))
$\begin{split}s^{\prime}&=S_{x}^{2}\,s,\end{split}$ (34)
$\begin{split}\Sigma^{\prime}&=A\,\Sigma\,A^{T},\end{split}$ (35)
$\begin{split}\tau^{\prime}&=S_{t}^{2}\,\tau,\end{split}$ (36)
$\begin{split}v^{\prime}&=\frac{S_{x}}{S_{t}}(A\,v+u).\end{split}$ (37)
By restricting this result to the purely spatial case, when temporal dependencies are disregarded, it holds that for purely spatial smoothing kernels of the form
$T(x;\;s,\Sigma)=g(x;\;s,\Sigma)$ (38)
the corresponding purely spatially smoothed image representations are related according to
$L^{\prime}(x^{\prime};\;s^{\prime},\Sigma^{\prime})=L(x;\;s,\Sigma),$ (39)
provided that the parameters of the purely spatial receptive fields are related according to (34) and (35).
In this way, these results imply that the essential components of the receptive fields in terms of either the purely spatial or the joint spatio-temporal smoothing transformations can be perfectly matched between the image data before and after the composed geometric image transformation. Thereby, these covariance properties provide a way of expressing an identity operation between the spatial or spatio-temporal smoothing effects of the idealized receptive field models, under compositions of (i) spatial scaling transformations, (ii) affine image transformations, (iii) Galilean transformations and (iv) temporal scaling transformations.
A consequence of these results is, however, that in order to make it possible to match the spatially or spatio-temporally smoothed image data between two observations of the same object under different viewing conditions, we have to expand the representation of the receptive field responses over multiple values of the parameters of the receptive fields; the set of parameters $(s,\Sigma)$ in the purely spatial case or the set of parameters $(s,\Sigma,v,\tau)$ in the joint spatio-temporal case. In other words, the shapes of the visual receptive fields should be expanded over the degrees of freedom of the class of geometric image transformations. We will return to that topic in Section 6 of this treatment.
5.2 Transformation properties of the pure spatial and temporal derivative operators
In addition to transforming the effect of the purely spatial or joint spatio-temporal smoothing operation in the idealized receptive field models (18) and (4.3), we do additionally have to consider how to transform the effects of the spatial and the temporal differentiation operators in the idealized receptive field models according to (18) and (4.3).
Formally, under the classes of primitive geometric image transformations in Equations (1)–(4), we have the following transformation properties for the purely spatial and temporal derivative operators:
Uniform spatial scaling transformations:
With $\nabla_{x}=(\partial_{x_{1}},\partial_{x_{2}})^{T}$ denoting the spatial gradient operator, spatial derivatives between the two image domains transform according to
$\nabla_{x^{\prime}}=\frac{1}{S_{x}}\,\nabla_{x},$ (40)
implying that directional derivatives over the two image domains defined according to
$\begin{split}\partial_{\varphi}=e_{\varphi}^{T}\,\nabla_{x},\end{split}$ (41)
$\begin{split}\partial_{\varphi^{\prime}}=e_{\varphi^{\prime}}^{T}\,\nabla_{x^{\prime}},\end{split}$ (42)
are related according to
$\partial_{\varphi^{\prime}}=\frac{1}{S_{x}}\,\partial_{\varphi}.$ (43)
Spatial affine transformations:
Spatial derivatives between the two image domains transform according to [^7] Concerning the notation, we throughout this paper denote the transpose of an inverse matrix as $A^{-T}=(A^{-1})^{T}$.
$\nabla_{x^{\prime}}=A^{-T}\,\nabla_{x}.$ (44)
Galilean transformations:
With $\nabla_{(x,t)}=(\partial_{x_{1}},\partial_{x_{2}},\partial_{t})^{T}$ denoting the spatio-temporal gradient operator and with the $3\times 3$ matrix $G$ representing the effect of the Galilean transformation (3) on the form
$\left(\begin{array}[]{c}x_{1}^{\prime}\\ x_{2}^{\prime}\\ t^{\prime}\end{array}\right)=G\left(\begin{array}[]{c}x_{1}\\ x_{2}\\ t\end{array}\right)=\left(\begin{array}[]{c}x_{1}-u_{1}\,t\\ x_{2}-u_{2}\,t\\ t\end{array}\right),$ (45)
spatio-temporal derivatives between the two image domains transform according to
$\nabla_{(x^{\prime},t^{\prime})}=G^{-T}\,\nabla_{(x,t)},$ (46)
where
$G=\left(\begin{array}[]{ccc}1&0&-u_{1}\\ 0&1&-u_{2}\\ 0&0&1\end{array}\right)$ (47)
and
$G^{-T}=\left(\begin{array}[]{ccc}1&0&0\\ 0&1&0\\ u_{1}&u_{2}&1\end{array}\right).$ (48)
Temporal scaling transformations:
Temporal derivatives between the two image domains transform according to
$\partial_{t^{\prime}}=\frac{1}{S_{t}}\,\partial_{t}.$ (49)
A fundamental limitation of using such pure spatial and temporal derivative operators in the idealized receptive field models, however, is that the magnitudes of the corresponding receptive field responses over the image domain after the geometric image transformation may be strongly different from the magnitudes of the receptive field responses over the image domain before the geometric image transformation. Thereby, it would be very hard to establish a direct matching between the receptive field responses before and after the geometric image transformation, based on the magnitudes of the receptive field responses, thus totally breaking the effect of the matching effects established by covariance properties of the purely spatial smoothing operation in (39) or the joint spatio-temporal smoothing operation in (33).
5.3 Individual covariance properties for idealized models of receptive fields based on scale-normalized spatial and temporal derivatives
A powerful way of avoiding the problem described in the previous section, that the magnitudes of spatial and temporal derivatives may be strongly influenced by the particular form of the geometric image transformation, is by instead introducing scale-normalized derivative operators, whose magnitudes can be perfectly matched under the influence of geometric image transformations.
5.3.1 Scale-normalized spatial derivatives
To handle the effect of uniform spatial scaling transformations on spatial image data, we can introduce scale-normalized spatial derivatives corresponding to the regular spatial gradient operator $\nabla_{x}=(\partial_{x_{1}},\partial_{x_{2}})^{T}$ according to (Lindeberg 1998 Equation (6)) (here simplified by setting the more general scale normalization parameter $\gamma$ to $\gamma=1$)
$\nabla_{x,\text{norm}}=s^{1/2}\,\nabla_{x},$ (50)
where $s=\sigma_{x}^{2}$ denotes the spatial scale parameter of the here assumed isotropic Gaussian kernel (with its covariance matrix being equal to a unit matrix $\Sigma=I$) used for performing the spatial smoothing. The corresponding scale-normalized directional derivative operator in the direction $e_{\varphi}=(\cos\varphi,\sin\varphi)^{T}$ then becomes
$\partial_{\varphi,\text{norm}}=s^{1/2}\,\partial_{\varphi}=s^{1/2}\,e^{T}\nabla_{x}.$ (51)
If we define the isotropic spatial scale-space representation $L\colon{\mathbb{R}}^{2}\times{\mathbb{R}}_{+}\rightarrow{\mathbb{R}}$ of any purely spatial image $f\colon{\mathbb{R}}^{2}\rightarrow{\mathbb{R}}$ according to
$L(\cdot;\;s)=g(\cdot;\;s,I)*f(\cdot),$ (52)
then it can be shown (Lindeberg 1998 Section 4.1; Lindeberg 2025d Sections 3.1–3.2) that under a uniform spatial scaling transformation of the form (1)
$f^{\prime}(x^{\prime})=f(x)\quad\quad\mbox{for}\quad\quad x^{\prime}=S_{x}\,x$ (53)
for matching values of the spatial scale parameters according to (34)
$s^{\prime}=S_{x}^{2}\,s,$ (54)
the corresponding scale-normalized derivatives for the transformed spatial scale-space representation
$L^{\prime}(\cdot;\;s)=g(\cdot;\;s^{\prime},I)*f^{\prime}(\cdot),$ (55)
will be equal at corresponding image points $x^{\prime}=S_{x}\,x$ according to
$\begin{split}(\nabla_{x^{\prime},\scriptsize\mbox{norm}}L^{\prime})(x^{\prime};\;s^{\prime})&=(\nabla_{x,\scriptsize\mbox{norm}}L)(x;\;s),\end{split}$ (56)
$\begin{split}(\nabla_{x^{\prime},\scriptsize\mbox{norm}}\nabla_{x^{\prime},\scriptsize\mbox{norm}}^{T}L^{\prime})(x^{\prime};\;s^{\prime})&=(\nabla_{x,\scriptsize\mbox{norm}}\nabla_{x,\scriptsize\mbox{norm}}^{T}L)(x;\;s),\end{split}$ (57)
$\begin{split}L^{\prime}_{{\varphi^{\prime}}^{m},\scriptsize\mbox{norm}}(x^{\prime};\;s^{\prime})&=L_{\varphi^{m},\scriptsize\mbox{norm}}(x;\;s).\end{split}$ (58)
Here, the first expression (56) represents (regular) scale-normalized gradient operators over the domains after and before the spatial scaling transformation. The second expression (57) represents (regular) scale-normalized Hessian operators $({\cal H}_{x^{\prime}}L^{\prime})(x^{\prime};\;s^{\prime})$ and $({\cal H}_{x}L)(x;\;s)$ over the domains after and before the image transformation. The third expression (58) represents (regular) scale-normalized directional derivative operators of order $m$ over the domains after and before the geometric transformation.
By this use of (regular) scale-normalized spatial derivatives, the spatial and the spatio-temporal receptive fields according to the generalized Gaussian derivative model will be provably covariant under uniform spatial scaling transformations of the form (1). In this way, a visual system built from such computational primitives will be able to handle objects of different size in the world as well as at different distances to the visual observer in a similar manner, as previously extensively explored to compute scale-covariant and scale-invariant image representations in the area of classical computer vision (Lindeberg 2021a).
5.3.2 Affine-normalized spatial derivatives
To handle the effect of more general spatial affine transformations
$f^{\prime}(x^{\prime})=f(x)\quad\quad\mbox{for}\quad\quad x^{\prime}=A\,x$ (59)
on spatial image data, one needs to make use of different spatial covariance matrices $\Sigma$ and $\Sigma^{\prime}$ for the spatial receptive fields before and after the spatial affine transformation. For this purpose, three main notions [^8] Notably, in the special case when the spatial covariance matrix $\Sigma$ in the affine Gaussian kernel is equal to a unit matrix $\Sigma=I$, the affine Gaussian kernel in (60) reduces to the isotropic Gaussian matrix in (52), in turn implying that the affine scale-normalized directional derivative operator in (70) reduces to the regular scale-normalized directional derivative operator used in (58). Furthermore, when $\Sigma=I$ the scale-normalized affine gradient operator in (70) similarly reduces to the regular scale-normalized gradient operator used in (56). Similarly, when $\Sigma=I$ the scale-normalized affine Hessian operator in (74) reduces to the regular scale-normalized Hessian operator used in (57). In these respects, the affine-normalized derivative operators in Section 5.3.2 constitute generalizations of the previously used regular (isotropic) scale-normalized derivative operators in Section 5.3.1 from an isotropic spatial scale space representation generated by convolution with rotationally symmetric Gaussian kernels to an anisotropic affine Gaussian scale space generated by convolution with anisotropic affine Gaussian kernels. of affine-normalized spatial derivatives have been proposed in Lindeberg (2025d) Sections 3.3–3.8.
The theoretical background to these affine-normalized spatial derivatives is that we consider the spatial affine scale-space representation $L\colon{\mathbb{R}}^{2}\times{\mathbb{R}}_{+}\times{\mathbb{S}}_{+}^{2}\rightarrow{\mathbb{R}}$ of any 2-D purely spatial image $f\colon{\mathbb{R}}^{2}\rightarrow{\mathbb{R}}$ defined according to
$L(\cdot;\;s,\Sigma)=g(\cdot;\;s,\Sigma)*f(\cdot),$ (60)
generated by convolution with affine Gaussian kernels $g\colon{\mathbb{R}}^{2}\times{\mathbb{R}}_{+}\times{\mathbb{S}}_{+}^{2}\rightarrow{\mathbb{R}}$ according to (26) and having spatial covariance matrices $\Sigma$ not necessarily equal to a unit matrix, with its corresponding transformed affine spatial scale-space representation
$L^{\prime}(\cdot;\;s^{\prime},\Sigma^{\prime})=g(\cdot;\;s^{\prime},\Sigma^{\prime})*f(\cdot),$ (61)
which obeys the affine covariance property (Lindeberg and Gårding 1997 Equation (29))
$L^{\prime}(x^{\prime};\;s^{\prime},\Sigma^{\prime})=L(x;\;s,\Sigma)$ (62)
for matching values of the scale parameters $s$ and $s^{\prime}$ and the spatial covariance matrices $\Sigma$ and $\Sigma^{\prime}$ according to (Lindeberg 2025d Equation (118))
$s^{\prime}\,\Sigma^{\prime}=s\,(S_{x}\,A)\,\Sigma\,(S_{x}A)^{T}=s\,S_{x}^{2}\,A\,\Sigma\,A^{T}.$ (63)
Then, given these affine Gaussian scale-space representations over the domains before and after the affine transformation, affine-normalized spatial derivative operators can be defined as follows:
The affine scale-normalized directional derivative:
A straightforward way of defining an extension of scale-normalized derivatives when using a spatial covariance matrix not equal to the unit matrix is according to (Lindeberg 2025d Equation (33))
$\partial_{\varphi,\scriptsize\mbox{norm}}^{m}=s^{m/2}\,(e_{\varphi}^{T}\,\Sigma\,e_{\varphi})^{m/2}\,\partial_{\varphi}^{m},$ (64)
where the entity $e_{\varphi}^{T}\,\Sigma\,e_{\varphi}$ reflects the amount of spatial smoothing in the direction $e_{\varphi}$, see Figure 9 for an illustration.
Figure 9: Illustration of the definition of the scale-normalization factor in the definition of the affine scale-normalized derivative operator according to (64). For performing the scale normalization in the direction $e_{\varphi}=(\cos\varphi,\sin\varphi)$ of the spatial covariance matrix $\Sigma$, an ellipse representation of the spatial covariance matrix $\Sigma$ is intersected in that direction, thus projecting the spatial smoothing effect of corresponding affine Gaussian kernel in the direction of the directional derivative operator. (Figure reproduced from Lindeberg (2025d) with permission (Open Access).)
In Lindeberg (2025d) Section 3.4, it is shown that this notion of affine scale-normalized directional derivatives is covariant under the similarity group, that is under combinations of uniform spatial scaling transformations and rotations. This notion of affine scale-normalized directional derivatives is also covariant in the special configuration when the affine transformation matrix $A$ and the spatial covariance matrix $\Sigma$ have the same eigenvectors, with the geometric interpretation that such a configuration corresponds to varying the viewing direction along the tilt [^9] The tilt direction is the projection of the local surface normal onto the image plane. direction of an observed local surface patch. Thus, for these subgroups of the group of spatial affine transformations, the affine scale-normalized directional derivatives will be equal over the domains before and after these special forms of spatial affine transformations:
$\partial_{\varphi^{\prime},\scriptsize\mbox{norm}}^{m}L^{\prime}(x^{\prime};\;s^{\prime},\Sigma^{\prime})=\partial_{\varphi,\scriptsize\mbox{norm}}^{m}L(x;\;s,\Sigma).$ (65)
As shown in Lindeberg (2025d) Section 3.4.5, the affine scale-normalized directional derivatives are, however, not covariant under fully general affine transformations.
Figure 10: Illustration of the covariance property (72) of the scale-normalized affine gradient operator according to (70) under general (non-singular) affine transformations. The interpretation of this covariance property is that, if we consider two cameras, that view the same local surface patch from general (non-degenerate) viewing conditions, then, to first order of approximation, the resulting affine gradient responses for the different views, here illustrated as arrows before the affine scale normalization, can, up to a rotation transformation $\tilde{\rho}$, be perfectly matched, provided that the scale parameters and the covariance matrices of the receptive fields are properly matched. (Figure reproduced from Lindeberg (2025d) with permission (Open Access).)
The scale-normalized affine gradient:
Given an eigenvalue decomposition of the $2\times 2$ symmetric and positive definite spatial covariance matrix $\Sigma$ of the form
$\Sigma=U\Lambda\,U^{T},$ (66)
where $\Lambda={\operatorname{diag}}(\lambda_{1},\lambda_{2})$ is a $2\times 2$ diagonal matrix with positive elements, and $U$ is a $2\times 2$ real unitary matrix, let us define the principal square root of $\Sigma$ as
$\Sigma^{1/2}=\Lambda^{1/2}\,U^{T},$ (67)
where $\Lambda^{1/2}={\operatorname{diag}}(\lambda_{1}^{1/2},\lambda_{2}^{1/2})$. Let us, however, note that the definition of the square root of a $2\times 2$ matrix is not unique, since for any arbitrary $2\times 2$ rotation matrix $\rho$, also the matrix
$\Sigma^{1/2}=\rho\,\Lambda^{1/2}\,U^{T},$ (68)
satisfies
$(\Sigma^{1/2})^{T}(\Sigma^{1/2})=U\,\Lambda^{1/2}\,\rho^{T}\,\rho\,\Lambda^{1/2}\,U^{T}=U\Lambda\,U^{T}.$ (69)
Given this definition of the principal root of the spatial covariance matrix $\Sigma$, we can define the scale-normalized affine gradient operator as (Lindeberg 2025d Equation (111))
$\nabla_{x,\scriptsize\mbox{affnorm}}=s^{1/2}\,\Sigma^{1/2}\,\nabla_{x}.$ (70)
Under a spatial affine transformation of the form (2), it is shown in Lindeberg (2025d) Section 3.6 that the scale-normalized affine gradient operator over the transformed domain $\nabla_{x^{\prime},\scriptsize\mbox{affnorm}}$ is up to a rotation matrix $\tilde{\rho}$ related to the scale-normalized affine gradient operator $\nabla_{x,\scriptsize\mbox{affnorm}}$ over the original domain according to (Lindeberg 2025d Equation (132))
$\nabla_{x^{\prime},\scriptsize\mbox{affnorm}}=\tilde{\rho}\,s^{1/2}\,\Sigma^{1/2}\,\nabla_{x}.$ (71)
Thereby, the scale-normalized affine gradient vectors $\nabla_{x,\scriptsize\mbox{affnorm}}L$ and $\nabla_{x^{\prime},\scriptsize\mbox{affnorm}}L^{\prime}$ computed from an affine scale-space representation of the form (60) over the domains before and after the affine transformation are related according to
$(\nabla_{x^{\prime},\scriptsize\mbox{affnorm}}L^{\prime})(x^{\prime};\;s^{\prime},\Sigma^{\prime})=\tilde{\rho}\,(\nabla_{x,\scriptsize\mbox{affnorm}}L)(x;\;s,\Sigma),$ (72)
see Figure 10 for an illustration.
In the special case when the affine transformation $A$ is in the similarity group, it is shown in Lindeberg (2024) Section 3.6 that the rotation matrix $\tilde{\rho}$ reduces to a unit matrix. In this special case, the scale-normalized affine gradients over the two domains before and after the image transformation are therefore guaranteed to be equal.
Figure 11: Illustration of the covariance property (75) of the scale-normalized affine Hessian operator according to (74) under general (non-singular) affine transformations. This covariance property means that, if we consider two cameras, that view the same local surface patch from general (non-degenerate) viewing conditions, then, to first order of approximation, the resulting affine Hessian responses for the different views, here illustrated as ellipses before the affine scale normalization, can, up to a combination of two (in this 2-D case related) rotation transformations $\tilde{\rho}$ and $\tilde{\rho}^{T}$, be perfectly matched, provided that the scale parameters and the covariance matrices of the receptive fields are properly matched. (Figure reproduced from Lindeberg (2025d) with permission (Open Access).)
The scale-normalized affine Hessian:
To extend the above notion from first- to second-order spatial derivatives, we can define a corresponding scale-normalized affine Hessian operator ${\cal H}_{x,\scriptsize\mbox{affnorm}}$ according to (Lindeberg 2025d Equation (140))
${\cal H}_{x,\scriptsize\mbox{affnorm}}=\nabla_{x,\scriptsize\mbox{affnorm}}\,\nabla^{T}_{x,\scriptsize\mbox{affnorm}},$ (73)
which, when expanded, then assumes the form
$\begin{split}{\cal H}_{x,\scriptsize\mbox{affnorm}}&=s\,(\Sigma^{1/2})\,\nabla_{x}\,\nabla_{x}^{T}(\Sigma^{1/2})^{T}.\end{split}$ (74)
Under a spatial affine transformation of the form (2), it can be shown that this operator transforms according to the following, between the domains before and after the image transformation (Lindeberg 2025d Equation (152))
${\cal H}_{x^{\prime},\scriptsize\mbox{affnorm}}=\tilde{\rho}\,{\cal H}_{x,\scriptsize\mbox{affnorm}}\,\tilde{\rho}^{T},$ (75)
where again $\tilde{\rho}$ denotes a rotation matrix. Thereby, the scale-normalized affine Hessian matrices computed from the affine Gaussian scale-space representations $L(x;\;s,\Sigma)$ and $L^{\prime}(x^{\prime};\;s^{\prime},\Sigma^{\prime})$ over the domains before and after the image transformation are related according to
$({\cal H}_{x^{\prime},\scriptsize\mbox{affnorm}}L^{\prime})(x^{\prime};\;s^{\prime},\Sigma^{\prime})=\\ =\tilde{\rho}\,({\cal H}_{x,\scriptsize\mbox{affnorm}}L)(x;\;s,\Sigma)\,\tilde{\rho}^{T},$ (76)
thus showing that the set of second-order spatial derivatives before and after the image transformation can up to an indeterminacy with respect to a possibly unknown rotation matrix be perfectly matched, see Figure 11 for an illustration.
Again, if the affine transformation matrix $A$ is in the similarity group, the rotation matrix $\rho$ reduces to a unit matrix.
In these ways, we can thus match receptive field responses formulated in terms of spatial derivatives of covariant spatial smoothing kernels under the spatial affine transformations that arise when viewing the same local surface patch from different viewing directions, with a few technical differences depending on the types of spatial derivative expression and the generality of the types of spatial affine transformations.
5.3.3 Scale-normalized temporal derivatives
To handle the effect of temporal scaling transformations, that model the effect of temporal or spatio-temporal events occurring either faster or slower relative to a reference view, one can introduce scale-normalized temporal derivatives according to Lindeberg (2017) Equation (6)
$\partial_{t,\scriptsize\mbox{norm}}^{n}=\tau^{n/2}\,\partial_{t}^{n},$ (77)
where $\tau=\sigma_{t}^{2}$ denotes the temporal scale parameter in the temporal scale-space representation $L\colon{\mathbb{R}}\times{\mathbb{R}}_{+}\rightarrow{\mathbb{R}}$ of a temporal signal $f\colon{\mathbb{R}}\rightarrow{\mathbb{R}}$ according to
$L(\cdot;\;\tau)=h(\cdot;\,\tau)*f(\cdot),$ (78)
with the temporal smoothing kernel $h\colon{\mathbb{R}}\times{\mathbb{R}}_{+}\rightarrow{\mathbb{R}}$ being either the non-causal 1-D Gaussian kernel according to (27) or the time-causal limit kernel according to (28).
With these definitions (and with the more general scale normalization power $\gamma$ in Lindeberg (2017) set to $\gamma=1$), the resulting scale-normalized derivatives of the transformed temporal scale-space representation
$L^{\prime}_{{t^{\prime}}^{n},\text{norm}}(t^{\prime};\;\tau^{\prime})={\tau^{\prime}}^{n/2}\,\partial_{t^{n}}(h(\cdot;\,\tau^{\prime})*f^{\prime}(\cdot))(t^{\prime};\;\tau^{\prime})$ (79)
do under a temporal scaling transformation of the form (4)
$f^{\prime}(t^{\prime})=f(t)\quad\quad\mbox{for}\quad\quad t^{\prime}=S_{t}\,t$ (80)
become equal at corresponding temporal moments $t^{\prime}=S_{t}\,t$ according to (Lindeberg 2017 Equation (10))
$L^{\prime}_{{t^{\prime}}^{n},\scriptsize\mbox{norm}}(t^{\prime};\;\tau^{\prime})=L_{t^{n},\scriptsize\mbox{norm}}(t;\;\tau)$ (81)
for matching values of the temporal scale parameters $\tau$ and $\tau^{\prime}$ over the domains before and after the temporal scaling transformation according to (36)
$\tau^{\prime}=S_{t}^{2}\,\tau.$ (82)
In this way, a vision system based on temporal filtering with scale-normalized temporal derivatives of either the non-causal Gaussian kernel or the time-causal limit kernel as the temporal smoothing kernel will be able to handle spatio-temporal events that occur either faster or slower between different views of an otherwise similar type of spatio-temporal event.
5.3.4 Scale-normalized velocity-adapted temporal derivatives
To handle the effect of Galilean transformations on spatio-temporal image data, one can extend the notion of scale-normalized temporal derivatives according to (77) into scale-normalized velocity-adapted temporal derivatives according to (Lindeberg 2025d Equation (168))
$\partial_{{\bar{t}},\scriptsize\mbox{norm}}^{n}=\tau^{n/2}\,(v^{T}\,\nabla_{x}+\partial_{t})^{n}.$ (83)
Then, under the simultaneous application of a Galilean transformation of the form (3) with potentially both spatial and temporal scaling transformations according to (1) and (4)
$\begin{split}x^{\prime}=S_{x}\,(x+u\,t),\end{split}$ (84)
$\begin{split}t^{\prime}=S_{t}\,t,\end{split}$ (85)
it holds that if we define a joint spatio-temporal scale-space representation $L\colon{\mathbb{R}}^{2}\times{\mathbb{R}}\times{\mathbb{R}}_{+}\times{\mathbb{R}}_{+}\times{\mathbb{R}}^{2}\rightarrow{\mathbb{R}}$ by convolving any video sequence or video stream $f\colon{\mathbb{R}}^{2}\times{\mathbb{R}}$ with the spatio-temporal smoothing kernel $T\colon{\mathbb{R}}^{2}\times{\mathbb{R}}\times{\mathbb{R}}_{+}\times{\mathbb{R}}_{+}\times{\mathbb{R}}^{2}\rightarrow{\mathbb{R}}$ according to
$T(x,t;\;s,\tau,v)=g(x-v\,t;\;s,I)\,h(t;\;\tau),$ (86)
then if we define the spatio-temporal scale-space representations $L$ and $L^{\prime}$ before and after the composed geometric image transformation
$f^{\prime}(x^{\prime},t^{\prime})=f(x,t)$ (87)
according to
$\begin{split}L(\cdot,\cdot;\,s,\tau,v)=T(\cdot,\cdot;\,s,\tau,v)*f(\cdot;\cdot),\end{split}$ (88)
$\begin{split}L^{\prime}(\cdot,\cdot;\,s^{\prime},\tau^{\prime},v^{\prime})=T(\cdot,\cdot;\,s^{\prime},\tau^{\prime},v^{\prime})*f^{\prime}(\cdot;\cdot),\end{split}$ (89)
the corresponding scale-normalized velocity-adapted temporal derivatives can be perfectly matched according to
$L^{\prime}_{{\bar{t^{\prime}}^{n}},\scriptsize\mbox{norm}}(x^{\prime},t^{\prime};\;s^{\prime},\tau^{\prime},v^{\prime})=L_{{\bar{t}}^{n},\scriptsize\mbox{norm}}(x,t;\;s,\tau,v),$ (90)
provided that the filter parameters over the domains before and after the composed geometric image transformation are matched according to
$\begin{split}s^{\prime}&=S_{x}^{2}\,s,\end{split}$ (91)
$\begin{split}\tau^{\prime}&=S_{t}^{2}\,\tau,\end{split}$ (92)
$\begin{split}v^{\prime}&=\frac{S_{x}}{S_{t}}(v+u).\end{split}$ (93)
In this way, a visual system based on spatio-temporal receptive fields that comprise velocity-adapted receptive based on such a velocity parameter $v$ in both the spatial smoothing kernel $T$ and the velocity-adapted temporal derivative operators $\partial_{{\bar{t}},\scriptsize\mbox{norm}}^{n}$ will have the ability to handle different types of relative motions, as parameterized by the velocity parameter $u$ between the viewing direction and the observer.
Note, however, that it is, in general, not sufficient to include a variability with respect to only the image velocity parameter $v$ in the model. Since the value of that parameter may be changed during the image transformation, also a variability is needed concerning the ratio between the spatial and the temporal scale parameters. Thereby, it is therefore necessary to also consider the potential interaction effects between the different types of primitive geometric image transformations (1)–(4) when modelling the combined effect of composed spatio-temporal image transformations.
5.4 Joint covariance properties for receptive fields in terms of spatial and spatio-temporal derivatives under the composed geometric transformation model
Let us next consider composed spatio-temporal image transformations according to (5) and (6) for the monocular projection model
$\begin{split}x^{\prime}=S_{x}\,(A\,x+u\,t),\end{split}$ (94)
$\begin{split}t^{\prime}=S_{t}\,t.\end{split}$ (95)
and according to (9) and (6) for the binocular projection model
$\begin{split}x^{\prime}=\tilde{B}\,x+\tilde{u}\,t,\end{split}$ (96)
$\begin{split}t^{\prime}=S_{t}\,t.\end{split}$ (97)
Then, it follows that:
  • If we define spatio-temporal receptive fields defined based on to the affine scale-normalized directional derivative operator according to (64)
    $\partial_{\varphi,\scriptsize\mbox{norm}}^{m}=s^{m/2}\,(e_{\varphi}^{T}\,\Sigma\,e_{\varphi})^{m/2}\,\partial_{\varphi}^{m}$ (98)
    and the scale-normalized velocity-adapted temporal derivative operator according to (83)
    $\partial_{{\bar{t}},\scriptsize\mbox{norm}}^{n}=\tau^{n/2}\,(v^{T}\,\nabla_{x}+\partial_{t})^{n},$ (99)
    then the composed spatio-temporal derivatives will for compositions of spatial transformations within the similarity group, for which the affine transformation matrix reduces to a rotation matrix $A=R_{\theta}$, Galilean transformations and temporal scaling transformations be equal at corresponding spatio-temporal image points
    $\partial_{{\varphi^{\prime}}^{m},\scriptsize\mbox{norm}}\,\partial_{{t^{\prime}}^{n},\scriptsize\mbox{norm}}^{m}L^{\prime}(x^{\prime},t^{\prime};\;s^{\prime},\Sigma^{\prime},\tau^{\prime},v^{\prime})=\\ =\partial_{\varphi^{m},\scriptsize\mbox{norm}}\,\partial_{t^{n},\scriptsize\mbox{norm}}^{m}L(x,t;\;s,\Sigma,\tau,v),$ (100)
    provided that the other parameters of the receptive fields are matched according to (34)–(37) such that (Lindeberg 2025d Equations (277)–(281))
    $\begin{split}s^{\prime}&=S_{x}^{2}\,s,\end{split}$ (101)
    $\begin{split}\varphi^{\prime}&=\varphi+\theta,\end{split}$ (102)
    $\begin{split}\Sigma^{\prime}&=R_{\theta}\,\Sigma\,R_{\theta}^{T},\end{split}$ (103)
    $\begin{split}\tau^{\prime}&=S_{t}^{2}\,\tau,\end{split}$ (104)
    $\begin{split}v^{\prime}&=\frac{S_{x}}{S_{t}}(R_{\theta}\,v+u).\end{split}$ (105)
  • If we consider the group of general affine transformation matrices $A$, and define the scale-normalized affine gradient vector according to (70)
    $\nabla_{x,\scriptsize\mbox{affnorm}}=s^{1/2}\,\Sigma^{1/2}\,\nabla_{x},$ (106)
    the scale-normalized affine Hessian operatore ${\cal H}_{x,\scriptsize\mbox{affnorm}}$ defined from the regular Hessian operator ${\cal H}_{x}=\nabla_{x}\nabla_{x}^{T}$ according to (74)
    ${\cal H}_{x,\scriptsize\mbox{affnorm}}=s\,(\Sigma^{1/2})\,{\cal H}_{x}\,(\Sigma^{1/2})^{T},$ (107)
    and the scale-normalized velocity-adapted temporal derivative operator according to (83)
    $\partial_{{\bar{t}},\scriptsize\mbox{norm}}^{n}=\tau^{n/2}\,(v^{T}\,\nabla_{x}+\partial_{t})^{n},$ (108)
    then under the composed geometric image transformation given by (94) and (95), the resulting composed spatio-temporal receptive field responses will be equal up to a rotation matrix $\tilde{\rho}$ according to
    $(\nabla_{x^{\prime},\scriptsize\mbox{affnorm}}\,\partial_{{t^{\prime}}^{n},\scriptsize\mbox{norm}}^{m}L^{\prime})(x^{\prime},t^{\prime};\;s^{\prime},\Sigma^{\prime},\tau^{\prime},v^{\prime})=\\ =\tilde{\rho}\,(\nabla_{x,\scriptsize\mbox{affnorm}}\,\partial_{t^{n},\scriptsize\mbox{norm}}^{m}L)(x,t;\;s,\Sigma,\tau,v)$ (109)
    and
    $({\cal H}_{x^{\prime},\scriptsize\mbox{affnorm}}\,\partial_{{t^{\prime}}^{n},\scriptsize\mbox{norm}}^{m}L^{\prime})(x^{\prime},t^{\prime};\;s^{\prime},\Sigma^{\prime},\tau^{\prime},v^{\prime})=\\ =\tilde{\rho}\,({\cal H}_{x,\scriptsize\mbox{affnorm}}\,\partial_{t^{n},\scriptsize\mbox{norm}}^{m}L)(x,t;\;s,\Sigma,\tau,v)\,\tilde{\rho}^{T},$ (110)
    provided that the scale parameters $s$ and $s^{\prime}$ as well as the spatial covariance matrices $\Sigma$ and $\Sigma^{\prime}$ are matched according to (Lindeberg 2025d Equation (118))
    $s^{\prime}\,\Sigma^{\prime}=s\,(S_{x}\,A)\,\Sigma\,(S_{x}A)^{T}=s\,S_{x}^{2}\,A\,\Sigma\,A^{T},$ (111)
    and provided that the other parameters of the receptive fields are matched according to (36)–(37)
    $\begin{split}\tau^{\prime}&=S_{t}^{2}\,\tau,\end{split}$ (112)
    $\begin{split}v^{\prime}&=\frac{S_{x}}{S_{t}}(A\,v+u).\end{split}$ (113)
  • Irrespective of any restrictions on the family of affine transformation matrices $A$, the velocity-adapted temporal derivative operators according to (83) will be equal (Lindeberg 2025d Equation (291))
    $\partial_{{\bar{t}}^{\prime},\scriptsize\mbox{norm}}^{n}L^{\prime}(x^{\prime},t^{\prime};\;s^{\prime},\Sigma^{\prime},\tau^{\prime},v^{\prime})=\\ =\partial_{{\bar{t}},\scriptsize\mbox{norm}}^{n}L(x,t;\;s,\Sigma,\tau,v),$ (114)
    provided that the parameters $s$, $s^{\prime}$, $\Sigma$, $\Sigma^{\prime}$, $\tau$, $\tau^{\prime}$, $v$ and $v^{\prime}$ of the receptive fields are matched according to Equations (34)–(37).
While the above results have been formulated based on the monocular projection model (94)–(95), corresponding results for the binocular projection model (96)–(97) can be obtained by setting the uniform spatial scaling factor to $S_{x}=1$ and then replacing the affine transformation matrix $A$ by the affine transformation matrix $\tilde{B}$ in Equations (109)–(114).
[…]
Figure 12: Commutative diagram for scale-normalized spatio-temporal derivative operators defined from the joint spatio-temporal receptive field model (115) under the composition of (i) a spatial scaling transformation, (ii) a spatial affine transformation, (iii) a Galilean transformation and (iv) a temporal scaling transformation according to (5) and (6). This commutative diagram, which should be read from the lower left corner to the upper right corner, means that irrespective of whether the input video sequence or video stream $f(x,t)$ is first subject to the composed transformation $x^{\prime}=S_{x}(A\,x+u\,t)$ and $t^{\prime}=S_{t}\,t$ and then filtered with a scale-normalized spatio-temporal derivative kernel $(\nabla_{x^{\prime},\scriptsize\mbox{affnorm}}\partial_{t^{\prime},\scriptsize\mbox{norm}}T)(x^{\prime},t^{\prime};\;s^{\prime},\Sigma^{\prime},\tau^{\prime},v^{\prime})$, or instead directly convolved with the scale-normalized spatio-temporal smoothing kernel $(\nabla_{x,\scriptsize\mbox{affnorm}}\partial_{t,\scriptsize\mbox{norm}}T)(x,t;\;s,\Sigma,\tau,v)$ and then subject to the same joint spatio-temporal transformation, we do then, up to a possibly unknown rotation transformation $\tilde{\rho}$, get the same result, provided that the parameters of the spatio-temporal smoothing kernels are related according to $s^{\prime}=S_{x}^{2}\,s$, $\Sigma^{\prime}=A\,\Sigma\,A^{T}$, $\tau^{\prime}=S_{t}^{2}\,\tau$ and $v^{\prime}=\frac{S_{x}}{S_{t}}(A\,v+u)$. (Adapted from Lindeberg (2025d) (Open Access).)
[…]
Figures 1313 illustrate these results in terms of commutative diagrams for spatio-temporal receptive field responses under geometric image transformations for the specific spatio-temporal receptive field model
$T_{x,\scriptsize\mbox{affnorm},{\bar{t}},\scriptsize\mbox{norm}}(x,t;\;s,\Sigma,\tau,v)=\\ =\nabla_{x,\scriptsize\mbox{affnorm}}\,\partial_{{\bar{t}},\scriptsize\mbox{norm}}\,T(x,t;\;s,\Sigma,\tau,v).$ (115)
Corresponding commutative diagrams can also be formulated for the other combinations of spatio-temporal receptive field operators with the general types of composed geometric image transformations.
In this way, we thus have a general framework for how spatio-temporal receptive field responses can be matched under compositions of (i) uniform spatial scaling transformations, (ii) spatial affine transformations, (iii) Galilean transformations and (iv) temporal scaling transformations.
Corresponding results for purely spatial receptive fields can in turn be obtained by fully removing all the explicit temporal dependencies from the above relationships, that is by removing all the occurrences of scale-normalized velocity-adapted temporal derivative operators $\partial_{{\bar{t}},\scriptsize\mbox{norm}}$ as well as removing all the explicit dependencies on time $t$, the temporal scale $\tau$, the temporal scaling factor $S_{t}$, as well as the velocity parameters $u$ and $v$.
5.5 Explicit examples of covariant receptive field families
Given the above theoretical results in the previous section, and stated more explicitly, these results thus mean that:
  • If purely spatial image data $f\colon{\mathbb{R}}^{2}\rightarrow{\mathbb{R}}$ are filtered with the family of spatial receptive fields
    $T_{\varphi^{m},\scriptsize\mbox{norm}}(x;\;s,\Sigma)=\partial_{\varphi,\scriptsize\mbox{norm}}^{m}\,g(x;\;s,\Sigma),$ (116)
    with the affine scale-normalized directional derivative operator $\partial_{\varphi,\scriptsize\mbox{norm}}^{m}$ according to (64), then the resulting spatial receptive field responses are covariant under the spatial similarity group, that is under combinations of spatial scaling transformations and spatial rotations.
  • If joint spatio-temporal image data $f\colon{\mathbb{R}}^{2}\times{\mathbb{R}}\rightarrow{\mathbb{R}}$ are filtered with the family of spatio-temporal receptive fields
    $T_{\varphi^{m},\scriptsize\mbox{norm},{\bar{t}}^{n},\scriptsize\mbox{norm}}(x,t;\;s,\Sigma,\tau,v)=\\ =\partial_{\varphi,\scriptsize\mbox{norm}}^{m}\,\partial_{{\bar{t}},\scriptsize\mbox{norm}}^{n}\,(g(x-vt;\;s,\Sigma)\,h(t;\;\tau)),$ (117)
    with the affine scale-normalized directional derivative operator $\partial_{\varphi,\scriptsize\mbox{norm}}^{m}$ according to (64) and the scale-normalized velocity-adapted temporal derivative operator $\partial_{{\bar{t}},\scriptsize\mbox{norm}}^{n}$ according to (83), then the resulting spatial receptive field responses are covariant under the spatial similarity group, combined with joint covariance properties under Galilean transformations and temporal scaling transformations.
  • If purely spatial image data $f\colon{\mathbb{R}}^{2}\rightarrow{\mathbb{R}}$ are filtered with the family of spatial receptive fields
    $T_{\nabla,\scriptsize\mbox{affnorm}}(x;\;s,\Sigma)=\nabla_{x,\scriptsize\mbox{affnorm}}\,g(x;\;s,\Sigma),$ (118)
    with the scale-normalized affine gradient operator $\nabla_{x,\scriptsize\mbox{affnorm}}$ according to (70), then the resulting spatial receptive field responses are covariant under both spatial scaling transformations and spatial affine transformations.
  • If joint spatio-temporal image data $f\colon{\mathbb{R}}^{2}\times{\mathbb{R}}\rightarrow{\mathbb{R}}$ are filtered with the family of spatio-temporal receptive fields
    $T_{\nabla,\scriptsize\mbox{affnorm},{\bar{t}}^{n},\scriptsize\mbox{norm}}(x,t;\;s,\Sigma,\tau,v)=\\ =\nabla_{x,\scriptsize\mbox{affnorm}}\,\partial_{{\bar{t}},\scriptsize\mbox{norm}}^{n}(g(x-vt;\;s,\Sigma)\,h(t;\;\tau)),$ (119)
    with the scale-normalized affine gradient operator $\nabla_{x,\scriptsize\mbox{affnorm}}$ according to (70) and the scale-normalized velocity-adapted temporal derivative operator $\partial_{{\bar{t}},\scriptsize\mbox{norm}}^{n}$ according to (83), then the resulting spatial receptive field responses are covariant under combinations of spatial scaling transformations, spatial affine transformations, Galilean transformations and temporal scaling transformations.
  • If purely spatial image data $f\colon{\mathbb{R}}^{2}\rightarrow{\mathbb{R}}$ are filtered with the family of spatial receptive fields
    $T_{{\cal H},\scriptsize\mbox{affnorm}}(x;\;s,\Sigma)={\cal H}_{x,\scriptsize\mbox{affnorm}}\,g(x;\;s,\Sigma),$ (120)
    with the scale-normalized affine Hessian operator ${\cal H}_{x,\scriptsize\mbox{affnorm}}$ according to (74), then the resulting spatial receptive field responses are covariant under both spatial scaling transformations and spatial affine transformations.
  • If joint spatio-temporal image data $f\colon{\mathbb{R}}^{2}\times{\mathbb{R}}\rightarrow{\mathbb{R}}$ are filtered with the family of spatio-temporal receptive fields
    $T_{{\cal H},\scriptsize\mbox{affnorm},{\bar{t}}^{n},\scriptsize\mbox{norm}}(x,t;\;s,\Sigma,\tau,v)=\\ ={\cal H}_{x,\scriptsize\mbox{affnorm}}\,\partial_{{\bar{t}},\scriptsize\mbox{norm}}^{n}(g(x-vt;\;s,\Sigma)\,h(t;\;\tau)),$ (121)
    with the scale-normalized affine Hessian operator ${\cal H}_{x,\scriptsize\mbox{affnorm}}$ according to (74) and the scale-normalized velocity-adapted temporal derivative operator $\partial_{{\bar{t}},\scriptsize\mbox{norm}}^{n}$ according to (83), then the resulting spatial receptive field responses are covariant under combinations of spatial scaling transformations, spatial affine transformations, Galilean transformations and temporal scaling transformations.
  • If joint spatio-temporal image data $f\colon{\mathbb{R}}^{2}\times{\mathbb{R}}\rightarrow{\mathbb{R}}$ are filtered with the family of spatio-temporal receptive fields
    $T_{{\bar{t}}^{n},\scriptsize\mbox{norm}}(x,t;\;s,\Sigma,\tau,v)=\\ =\partial_{{\bar{t}},\scriptsize\mbox{norm}}^{n}(g(x-vt;\;s,\Sigma)\,h(t;\;\tau)),$ (122)
    with the scale-normalized velocity-adapted temporal derivative operator $\partial_{{\bar{t}},\scriptsize\mbox{norm}}^{n}$ according to (83), then the resulting spatial receptive field responses are covariant under combinations of spatial scaling transformations, spatial affine transformations, Galilean transformations and temporal scaling transformations.
Notably, these theoretical results comprise combinations of spatial derivatives up to order 2 with temporal derivatives for any order of temporal differentiation. In these ways, we can thus formulate a rich set of both purely spatial and joint spatio-temporal receptive field models, that are provably covariant under combinations of 4 main types of primitive geometric image transformations according to Equations (1)–(4), as summarized in the composed geometric image transformations according to (5), (6) and (9).
5.6 Relationships to the influence of illumination variations
Regarding variabilities in image data caused by natural image transformations, we do in this treatment focus on the influence due to geometric image transformations. Regarding the influence of illumination variations, which also constitute a large source to variability in image data, it is, however, interesting to note that according to the theory in Lindeberg (2013a) Section 2.3 condensely summarized in Lindeberg (2021b) Section 3.4, it holds that if the image data $f$ used as input to the receptive fields are parameterized in terms of the logarithm of the intensities in the dimension of the incoming energy $\log I(x,y)$ or $\log I(x,y,t)$ according to
$f(x,y)\sim\log I(x,y)\quad\quad\mbox{or}\quad\quad f(x,y,t)\sim\log I(x,y,t),$ (123)
then the computed spatial or spatio-temporal receptive field responses in terms of either spatial derivatives, temporal derivatives or both will be automatically invariant under local multiplicative intensity transformations of the form
$I(x,y)\mapsto C\,I(x,y)$ (124)
or
$I(x,y,t)\mapsto C\,I(x,y,t),$ (125)
for any strictly positive local multiplication factor $C\in{\mathbb{R}}_{+}$. A similar invariance result holds concerning the influence of global exposure compensation mechanisms of a similar multiplicative form.
In this way, a substantial component of the influence due to illumination variations and exposure compensation mechanism can be directly handled in a straightforward manner.
In this context, it is interesting to note that the retinex theory of early vision (Land 1974; 1986) also makes use of a logarithmic brightness scale. An exposure mechanism on the retina that adapts the diameter of the pupil and the sensitivity of the photopigments in such a way that the relative range in the variability of the signal divided by the mean illumination is held constant, can also be seen as implementing an approximation of the derivative of a logarithmic transformation
$d(\log z)=\frac{dz}{z},$ (126)
see e.g. Peli (1990). Furthermore, in the area of psychophysics, the Weber-Fechner law states that the ratio
$\frac{\Delta I}{I}=k,$ (127)
between the threshold $\Delta I$ corresponding to a just noticeable difference in image intensity and the background intensity $I$ is constant over large ranges of magnitude variations, see e.g. Palmer (1999) pages 671–672, thus providing further support for the relevance of a logarithmic brightness scale.
Figure 14: Conceptual illustration of how sets of spatial and/or spatio-temporal receptive field responses computed over different values of the filter parameters $(s,\Sigma,v,\tau)$ of the receptive fields can be matched between different views of the same scene or spatio-temporal event under different viewing conditions, by making use of the matching relations in Equations (34)–(37) between the filter parameters for the two image domains before and after the geometric image transformation, under variabilities of the parameters $(S_{x},A,v,S_{t})$ or $(B,v,S_{t})$ of the primitive geometric image transformations in Equations (1)–(4), as combined into composed geometric image transformations according to either Equation (5) or Equation (9) in combination with Equation (6). Such matching of the receptive field responses under geometric image transformations is possible for receptive fields that obey provable covariance properties as exemplified in Section $5.$ (The arrows between the boxes indicate the information flow from the image acquisition stage to the matching stage. Regarding the visualized receptive fields in the middle box, for the spatial receptive fields in the left column, the coordinates are the purely spatial coordinates $(x_{1},x_{2})\in{\mathbb{R}}^{2}$, whereas for the spatio-temporal receptive fields, only one of the spatial coordinates is shown, thus with the spatio-temporal image coordinates $(x_{1},t)\in{\mathbb{R}}\times{\mathbb{R}}$. The receptive fields in the top parts of the middle box are separable over image space or joint space-time and are based on isotropic Gaussian smoothing over the spatial domain with the spatial covariance matrix $\Sigma$ equal to a unit matrix $I$. The purely spatial receptive fields in the left bottom part of the middle box are based on spatial smoothing with non-isotropic affine Gaussian kernels for which the spatial covariance matrix is not equal to a unit matrix. The joint spatio-temporal receptive fields in the right bottom part of the middle box are based on velocity-adapted spatio-temporal smoothing kernels in combination with velocity-adapted temporal derivatives. All the spatio-temporal receptive fields in this figure are based on temporal smoothing with the time-causal limit kernel (22).)
6 Do the shapes of the simple cells in the primary visual cortex of higher mammals span the variabilities of geometric image transformations to support explicitly covariant families of visual receptive fields?
A main result of the above presented theory is that the output from both purely spatial and joint spatio-temporal receptive fields according to generalized versions of the idealized receptive field models (18) and (4.3) can be matched under the composed geometric image transformations according to both the monocular projection model in Equations (94)–(95) and the binocular projection model in Equations (96)–(97). This result holds provided that we allow for the sets of the parameters $(s,\Sigma,\tau,v)$ and $(s^{\prime},\Sigma^{\prime},\tau^{\prime},v^{\prime})$ of the receptive fields $T(s,\Sigma,\tau,v)$ and $T^{\prime}(s^{\prime},\Sigma^{\prime},\tau^{\prime},v^{\prime})$ to be varied between the image domains before and after the geometric image transformation, and specifically having the values of the receptive field parameters being matched according to Equations (34)–(37) as functions of the parameters $(S_{t},A,u,S_{t})$ of the composed geometric image transformation.
Since the parameters $(S_{t},A,u,S_{t})$ of the geometric image transformation cannot be expected to be a priori known to a vision system that is to analyze an a priori unknown scene, a general purpose strategy for a vision system could therefore be to expand the receptive fields over a rich set of receptive fields, with the shapes of the receptive fields expanded over the degrees of freedom of the corresponding image transformations. Thereby, it would be possible to match the outputs from populations of receptive fields, to establish a corresponding matching of the receptive field responses obtained from a particular viewing condition in relation to a learned memory of receptive field responses computed from similar objects and spatio-temporal events under different sets of viewing conditions, see Figure 14 for an illustration. [^10] Note also that in an actual implementation of a matching scheme of receptive field responses of this form, one could also conceive that such a matching could be performed based on receptive field responses at higher levels in the visual hierarchy, which would indeed be possible based on the presented theory, if the receptive field responses at the higher levels are computed from the receptive field responses from the responses of the simple cells in a causal feed-forward manner that respects the covariance properties.
Given this idealized theory, one may therefore ask if biological vision has evolved similar mechanisms to be able to handle the influence of geometric image transformations on the receptive field responses in a way that is closely related to the results from the presented theory. Specifically, one may ask if the shapes of the receptive fields of the simple cells in the primary visual cortex are expanded over the degrees of freedom of (i) uniform spatial scaling transformations, (ii) non-isotropic spatial affine transformations, (iii) Galilean transformations and (iv) temporal scaling transformations, and if so for what species?
Given evolutionary arguments, one could argue that for at least higher species with more developed general purpose vision systems for addressing a wide range of visual tasks, like primates and cats, it would constitute an evolutionary advantage to adapt their visual system to the structure of the external environment, and thus to the influence of geometric image transformations on the image data that reach their visual sensors. One may therefore ask if there for such higher species under a sufficient evolutionary pressure would be variabilities in the shapes of their visual receptive fields in the visual area corresponding to V1, as implied by what would be the result of expanding the receptive field shapes with respect to the degrees of freedom of the 4 main classes of primitive geometric image transformations.
6.1 Purely spatial variabilities in the shapes of spatial and spatio-temporal receptive fields
In Lindeberg (2025a), the theoretical problem regarding variabilities in receptive field shapes as implied by covariance properties under geometric image transformations is addressed in detail in relation to the first degrees of freedom concerning the combined effect of (i) uniform spatial scaling transformations and (ii) non-isotropic spatial affine transformations. In brief, and extended from a purely spatial domain to also encompass the joint spatio-temporal domain, the results from that treatment are that:
Variability under uniform spatial scaling transformations:
Regarding the degree of freedom corresponding to uniform spatial scaling transformations, the corresponding degree of freedom in terms of the spatial scale parameter $\sigma_{x}=\sqrt{s}$ is special in the sense that the spatial affine Gaussian kernel obeys a semi-group property over spatial scales (see Lindeberg (1993) Equation (15.35))
$g(\cdot;\;\Sigma_{1})*g(\cdot;\;\Sigma_{2})=g(\cdot;\;\Sigma_{1}+\Sigma_{2}),$ (128)
from which it follows that any receptive field response at a coarser spatial scale, here as represented by the spatial gradient vector $\nabla_{x}L$ or the Hessian matrix ${\cal H}_{x}L$ for the spatial covariance matrix $\Sigma_{2}$, can be computed by affine Gaussian smoothing of the receptive field responses for any finer spatial scale, here as represented by the spatial gradient vector $\nabla_{x}L$ or the Hessian matrix ${\cal H}_{x}L$ defined for the spatial covariance matrix $\Sigma_{1}$
$\begin{split}(\nabla_{x}\,L)(\cdot;\;\Sigma_{2})=g(\cdot;\;\Sigma_{2}-\Sigma_{1})*(\nabla_{x}\,L)(\cdot;\;\Sigma_{1}),\end{split}$ (129)
$\begin{split}({\cal H}_{x}\,L)(\cdot;\;\Sigma_{2})=g(\cdot;\;\Sigma_{2}-\Sigma_{1})*({\cal H}_{x}\,L)(\cdot;\;\Sigma_{1}),\end{split}$ (130)
provided that the difference between the spatial covariance matrices
$\Delta\Sigma=\Sigma_{2}-\Sigma_{1}$ (131)
is a symmetric positive definite matrix.
Thereby, a vision system could in principle choose to compute the earliest layers of spatial receptive fields at only the finest spatial scale and nevertheless be able to compute the coarser spatial receptive fields at higher layers in the visual hierarchy. Thus, irrespective of whether the spatial receptive fields corresponding to the simple cells are expanded over the spatial scales, it seems very plausible that the vision system should have the ability to compute visual operations corresponding to spatial scale covariance.
Figure 4 shows an example of such a variability under spatial scaling variations for first-order spatial directional derivatives computed based on isotropic Gaussian smoothing. Figure 5 in Lindeberg (2025a) shows an example of such a variability extended to second-order spatial directional derivatives.
Variability under spatial rotations in the image plane:
From the structure of orientation maps in the primary visual cortex of higher mammals, as studied by Bonhoeffer and Grinvald (1991), Blasdel (1992), Koch et al. (2016) and others, it is clear that we can interpret these orientation maps as an expansion of the receptive field shapes over the image orientations, corresponding to the parameter $\varphi$ in the idealized receptive field models (18) and (4.3). Thereby, we could regard the vision system of higher mammals to have the ability to compute visual operations corresponding to rotation covariance.
Figure 15: Orientation map in the primary visual cortex of cat, as recorded by Koch et al. (2016) (OpenAccess), and and demonstrating that the visual cortex of higher mammals can be regarded as performing an explicit expansion of the receptive field shapes over spatial image orientations, as would be the result of combining the notion of covariance over the subgroup of pure image rotations.
Combined with spatial scale covariance, this means that we could regard the vision system of higher mammals to have the ability to be covariant under spatial similarity transformations, that is to combinations of spatial scaling transformations and spatial rotations.
Figure 6 in Lindeberg (2025a) shows examples of such a variability under spatial rotations for first- and second-order spatial directional derivatives computed based on non-isotropic affine Gaussian smoothing.
Variability under the degree of elongation of the receptive fields:
From studies of the orientation selectivity properties of simple cells established from neurophysiological cell recordings by Nauhaus et al. (2008) and Goris et al. (2015), it is clear that the simple and complex cells of primates and cat have a substantial variability in orientation selectivity properties, ranging from narrow to wide orientation selectivity properties, as illustrated in Figure 16.
Figure 16: Schematic depiction of results from measurements of the orientation tuning of neurons, at different positions in the visual cortex, adapted from Nauhaus et al. (2008), showing how the orientation tuning changes from broad to sharp, and thus higher degree of orientation selectivity, with increasing distance from the pinwheels, consistent with the qualitative behaviour that would be obtained if the ratio $\kappa$, between the scale parameters in underlying affine Gaussian smoothing step in the idealized models of spatial and spatio-temporal receptive fields, would increase when moving away from the centers of the pinwheels on the cortical surface. (From Lindeberg (2025c) with permission (OpenAccess).)
First-order simple cells Second-order simple cells Complex cells
Refer to caption Refer to caption Refer to caption
Figure 17: Graphs of the orientation selectivity for the idealized models of (top left) simple cells in terms of first-order directional derivatives of affine Gaussian kernels, (top middle) simple cells in terms of second-order directional derivatives of affine Gaussian kernels, (top right) complex cells in terms of directional quasi-quadrature measures that combine the first- and second-order simple cell responses in a Euclidean way. The parameter $\kappa$ represents the degree of elongation of the receptive fields, and corresponds to the ratio between the square roots of the eigenvalues of the covariance matrix $\Sigma$ of the affine Gaussian derivative-based receptive fields. (Horizontal axes: orientation $\theta\in[-\pi/2,\pi/2]$. Vertical axes: Amplitude of the receptive field response relative to the maximum response obtained for $\theta=0$.) (Adapted from Lindeberg (2025c) with permission (OpenAccess).)
From a theoretical analysis of the orientation selectivity properties of the affine Gaussian derivative and affine Gabor models of visual receptive fields in Lindeberg (2025b), we have established a connection that the degree of orientation selectivity increases with the degree of elongation of the receptive fields, see Figure 17. For the affine Gaussian derivative and affine Gabor models, that degree of elongation corresponds to the ratio between the eigenvalues of the affine Gaussian kernel used in these idealized models.
In Lindeberg (2025c), we have combined these two sources of biological and theoretical knowledge to propose that these results are consistent with the receptive fields of primates and cats spanning a variability in the degree of elongation of the receptive fields. In Lindeberg (2025a), we further show that this degree of freedom of the receptive fields corresponding to the degree of freedom spanned by the ratio between the singular values obtained from a singular value decomposition of the affine transformation matrix $A$.
Figure 7 in Lindeberg (2025a) shows examples of such a variability over the degree of elongation for first- and second-order spatial directional derivative operators computed based on non-isotropic affine Gaussian smoothing. Figure 5 in this paper shows an example of a combined variability over the degree of elongation of the receptive fields with spatial rotations in the image plane for first-order spatial directional derivatives based on non-isotropic affine Gaussian smoothing.
Variability under a 4:th purely spatial degree of freedom:
To span all the 4 degrees of freedom corresponding to the combination of uniform spatial scaling transformations with non-isotropic spatial affine transformations, there is one additional degree of freedom that corresponds to generalizing the idealized receptive field models (18) and (4.3) to not necessarily having the orientation $\varphi$ for computing the directional derivatives $\partial_{\varphi}^{m}$ being parallel to any of the eigendirections of the spatial covariance matrix $\Sigma_{\varphi}$.
As argued in in Lindeberg (2025a) Section 7.4, there have been exampels of biological receptive fields recorded by Yazdanbakhsh and Livingstone (2006) (see Figure 6 in that paper) that appear to be more similar to first- or second-order directional derivatives of Gaussian kernels in directions different from the principal directions of an affine Gaussian kernel compared to directional derivatives of such kernels in directions that coincide with the principal directions of affine Gaussian kernels.
Figure 8 in Lindeberg (2025a) shows an example of such a variability over the angle between the orientation for computing spatial directional derivatives relative to the principal eigendirections of the affine Gaussian kernel used for spatial smoothing.
In these ways, there is potential support in different respects for the hypothesis that the receptive fields of simple cells in the primary visual cortex of higher mammals ought to have the ability to be covariant under the combination of uniform spatial scaling transformations, rotations in the image plane and non-isotropic spatial affine transformations.
6.2 Additional variabilities involving the temporal domain
Additionally, by extending the arguments in Lindeberg (2023b):
Variability under Galilean transformations:
From the ability of the simple cells in the visual system of higher mammals to compute spatio-temporal receptive fields similar to velocity-adapted temporal derivatives (DeAngelis et al. 1995; 2004; Lindeberg 2021b Figure 18 bottom part) and the receptive fields in the area MT being able to compute velocity-dependent responses, it seems plausible that the visual system should be able to compute Galilean-covariant receptive field responses.
Figure 6 in this paper shows an example of such a variability under Galilean transformations for spatio-temporal receptive fields over a 1+1-D spatio-temporal domain, based on a first-order spatial derivative of a Gaussian kernel and a first-order temporal derivative of the time-causal limit kernel.
Variability under temporal scaling transformations:
Concerning the degree of freedom corresponding to temporal scaling transformations, the corresponding degree of freedom in terms of the temporal scale parameter $\sigma_{t}=\sqrt{\tau}$ is also special in the sense that both the non-causal temporal Gaussian kernel and time-causal limit kernel used for temporal smoothing in the idealized model (4.3) for spatio-temporal receptive fields obey cascade properties over temporal scales of the form (Lindeberg 2023a Equation (2))
$h(\cdot;\;\tau_{2})=(\Delta h)(\cdot;\;\tau_{1}\mapsto\tau_{2})*h(\cdot;\;\tau_{1})$ (132)
for any pair of temporal scales $(\tau_{1},\tau_{2})$ with $\tau_{2}>\tau_{1}$ and for some family of transformation kernels $(\Delta h)(t;\;\tau_{1}\mapsto\tau_{2})$. This implies that any receptive field responses, here represented as the temporal derivative $L_{t^{n}}(\cdot;\;\tau)$ at any coarser temporal scale $\tau_{2}$ can be computed by an additional temporal filtering operation being applied to the receptive field responses, here represented as the temporal derivatives $L_{t^{n}}(\cdot;\;\tau)$ at any finer temporal scale $\tau_{1}$, according to
$L_{t^{n}}(\cdot;\;\tau_{2})=(\Delta h)(\cdot;\;\tau_{1}\mapsto\tau_{2})*L_{t^{n}}(\cdot;\;\tau_{1}).$ (133)
In the case of a non-causal temporal domain, the difference kernel $\Delta h$ is a Gaussian kernel $(\Delta h)(t;\;\tau_{1}\mapsto\tau_{2})=g(t;\;\tau_{2}-\tau_{1})$ because of the semi-group property of the Gaussian kernel. In the case of a time-causal domain, the difference kernel is either a single truncated exponential kernel $h_{\exp}(t;\;\mu_{k})=1/\mu_{k}\exp(-t/\mu_{k})$ if $t>0$, or 0 otherwise, or the convolution of a set of such truncated exponential kernels, with appropriate time constants $\mu_{k}$ to fit the relationship between the temporal scale levels $\tau_{1}$ and $\tau_{2}$.
This, in turn, means that a vision system could, in principle, choose to only implement the earliest layers of temporal receptive fields at the finest temporal scale and nevertheless have the ability to compute the representations at coarser temporal scales, based on additional temporal smoothing applied to the temporal or spatio-temporal receptive field representations at the finest temporal scales. Thus, irrespective of whether the temporal receptive fields are expanded over the temporal scales, it seems plausible that the vision system should have the ability to compute visual operations corresponding to temporal scale covariance.
Figure 7 shows an example of such a variability under temporal scaling transformations for the purely temporal time-causal limit kernel.
An extended theory for how an idealized vision system, based on receptive fields in terms of the generalized Gaussian derivative theory for visual receptive fields, could compute receptive field responses at coarser spatial and temporal scales, based on cascade smoothing properties from finer spatial and temporal scales, is given in Lindeberg (2025g).
6.3 Outlines to further research to characterize variabilities of visual receptive fields with regard to variabilities in relation to geometric image transformations
Given the above theoretically motivated predictions, main questions for further research concern:
  • For which of the above variabilities can neurophysiological support be found regarding expansions of the shapes of the receptive fields over the corresponding degrees of freedom for different types of species?
  • Are the degrees of freedom corresponding to spatial scaling transformations and temporal scaling transformations special, because of the properties that the smoothing operations over the spatial and the temporal scale parameters can be performed in terms of cascade smoothing operations, implying that only representations of explicit receptive field responses at the finest spatial and/or temporal scales would be needed to compute the receptive field responses at coarser and/or temporal scales?
    Does efficiency call for the visual system to only implement the finest spatial and temporal scales in the sensorium, or does the vision system implement redundancies in this respect?
  • If there would be expansions over both the degrees of freedom of the affine group beyond uniform scaling transformations and rotations as well as an expansion over the degrees of freedom of the Galilean group, are those expansions then performed jointly or separately? In other words, are such expansions performed over the tensor product of the corresponding affine and Galilean parameters or separately for each subset of these degrees of freedom?
    For efficiency reasons, separate expansions over the affine and the Galilean degrees of freedom would need much less wetware to span the degrees of freedom of the corresponding geometric image transformations. A joint expansion over these degrees of freedom could on the other hand allow for the computation of more accurate estimates of the structure of dynamic scenes.
  • How well do the idealized models for spatial and spatio-temporal receptive fields explicitly listed in Section 5.5 model the biological receptive fields for simple cells in the primary visual cortex for different species?
  • How wide ranges in the parameter spaces of the receptive fields are spanned for the respective degrees of freedom for different species?
Answering these questions could provide valuable cues about how well different types of species have adapted their visual systems to handle the influence of different types of geometric image transformations on the image data.
Unfortunately, previous neurophysiological studies of visual receptive fields may not have been designed to answer the questions raised from this theoretical background. Answering the question about variabilities in the shapes of the visual receptive fields corresponding to the simple cells in the primary cortex would call for the accumulation of shape parameters of reconstructed receptive fields for a relatively large number of visual neurons, as well as using visual stimuli that would probe against the listed types of variabilities with regard to the 4 main types of geometric image transformations as leading to a total number of 7 degrees of freedom.
Currently, it seems very hard to get access to raw reconstructed receptive field profiles from previous recordings of the visual receptive fields in V1 of higher mammals. Publicly available receptive fields, as published in scientific publications, typically only show a rather small number of visual receptive fields. Concerning the hypothesized variabilities with regard to the degree of elongation of visual receptive fields, the neurophysiological evidence, that we have based our previous hypothesis on, are based on orientation selectivity properties of higher mammals (primates or cats) with a clear pinwheel structure. Recordings from mice, for which there is the largest number of public data, may on the other hand not be able to fully answer this question, because the vision system of mice is far less developed compared the visual systems of higher mammals (see e.g. Huberman and Niell (2011)), and does specifically lack pinwheel structures. In addition, recent results by Fu et al. (2024) of reconstructing most exciting inputs (MEIs) for early visual receptive fields in mice suggest that complex spatial features emerge earlier in the visual pathway of mice compared to primates, and that the receptive fields of mice are therefore more complex than a Gabor or Gaussian derivative model, as used in the theoretical model for stating the predictions in this paper.
For these overall reasons, in Lindeberg (2023b) Sections 3.2.1–3.2.2 and Lindeberg (2025c) Sections 4.2–4.3, sets of suggestions for more specific neurophysiological experiments are proposed to investigate the underlying hypotheses in more detail, and to characterize the structure of biological receptive fields in the primary visual cortex with respect to the influence of parameters of the receptive fields corresponding to the different degrees of freedom of the 4 main types of geometric image transformations. Complementary suggestions for further research in relation to the influence of geometric image transformations on early vision are also outlined in Lindeberg (2025a) Section 8.1.
Additionally, in Lindeberg (2025e) it is shown that that the results concerning an expansion over the degree of elongation for simple cells appear to extend to complex cells.
7 Summary and discussion
We have presented a principled theory for the interaction between geometric image transformations and receptive field responses, and used results from that theory to address the question about variabilities in the shapes of the receptive fields of simple cells in the primary visual cortex.
This theory is based on idealized models of visual receptive fields in terms of combinations of smoothing with spatial smoothing kernels of the form (26) or spatio-temporal smoothing kernels of the form (25) with scale-normalized spatial and temporal derivatives according to Section 5.3. In Sections 5.4 and 5.5, we have described how the resulting idealized models of spatial or spatio-temporal receptive fields obey provable covariance properties under compositions of spatial scaling transformations, spatial affine transformations, Galilean transformations and temporal scaling transformations. Specifically, we have in Section 6 considered the hypothesis about whether the receptive fields of simple cells in the primary visual cortex can be regarded as having their shapes expanded with regard to the degrees of freedom of the basic types of geometric image transformations that occur in the image formation process.
By postulating that the responses of idealized models of receptive fields in terms of scale-normalized spatial and temporal derivative operators should be possible to match between the image domains before and after the geometric image transformations, we have predicted a set of variabilities over (i) spatial scaling transformations, (ii) image rotations, (iii) the degree of spatial elongation of the receptive fields, (iv) a 4:th spatial degree of freedom, (v) Galilean transformation over image space-time and (vi) temporal scaling transformations. We have considered potential support for the covariance properties of the either purely spatial or joint spatio-temporal receptive fields with regard to these 7 degrees of freedom (the Galilean transformation comprises 2 degrees of freedom), in view of neurophysiological evidence and structural properties regarding how populations of receptive field responses can be computed based on structural properties of the families of receptive fields under variabilities over the filter parameters. In the absence of sufficient neurophysiological or psychophysical evidence to firmly state whether the predicted properties would hold in the primary visual cortex of higher mammals, we have in Section 6.3 pointed to directions for future neurophysiological investigations to characterize these properties in more detail.
Concerning a possible expansion of the shapes of the simple cells with regard to the degrees of freedom of the considered 4 main types of geometric image transformations in terms of (i) spatial scaling transformations, (ii) affine image transformations, (iii) Galilean transformations and (iv) temporal scaling transformations, it is of interest to consider the number of receptive fields in the early layers of the visual hierarchy. Given that the 1 M output channels from the retina are mapped to 1 M output channels from the lateral geniculate nucleus (LGN) to the primary visual cortex (V1) to 190 M neurons in V1 with 37 M output channels (see DiCarlo et al. (2012) Figure 3), the substantial expansion of the number of receptive fields from the LGN to V1 would indeed be consistent with an expansion of the shapes of the receptive fields over shape parameters of the receptive fields.
Furthermore, we can physically interpret the parameters $(S_{x},A,v,S_{t})$ of the primitive geometric image transformations in the composed geometric image transformations according to (5), (6) and (9) as follows (Lindeberg 2025d Section 9):
  • the spatial scaling factor $S_{t}$ corresponds to the inverse depth $1/Z$, if the affine transformation matrix $A$ in the composed monocular image transformation (5) is normalized in such a way that the affine transformation matrix $A$ reflects a scaled orthographic projection model,
  • knowledge about the affine transformation matrix $A$ in the composed monocular transformation model (5) provides direct information about the local surface orientation of the viewed local surface patch, according to the theoretical analysis in Gårding and Lindeberg (1996) Section 5.2,
  • knowledge about the affine transformation $B$ in the binocular transformation model (9) provides direct information about the local surface orientation of the viewed local surface patch, according to the theoretical analysis in Gårding and Lindeberg (1996) Section 6.1,
  • knowledge about image velocity $u$ in the monocular projection model (5), in combination with an estimate of the local depth $Z$ according to above, reveals the projection of the 3-D motion vector $U$ of the viewed object onto the image plane.
Hence, a vision system that is able to extract these parameters of the geometric image transformations based on processing and comparing populations of receptive field responses, should in principle have the ability to compute direct cues to the structure of environment, directly from established matching relations over the receptive field responses between the different views, or in relation to a learned memory of receptive field responses from previous views. Thereby, certain functionalities of the vision system could be formulated directly in terms of the parameters of the primitive geometric image transformations between different views of the same scene or the same spatio-temporal event.
Thus, in summary, we have presented a coherent theory for how the structure of the geometric image transformations, that determine the formation of the image data that reach the visual sensors in the perception system, influence the desired structure of the receptive field responses in an idealized vision system. Specifically, we have argued how we based on the requirement of formulating identity relations between receptive field responses acquired under different viewing conditions for the same scene, are lead to the requirement of covariance properties of the receptive fields, which in turn imply a set of variabilities of the shapes of the receptive fields, to span the degrees of freedom of the corresponding geometric image transformations. In this way, we have formulated a set of theoretical predictions, and also discussed partial support of these predictions in view of existing neurophysiological evidence and complementary theoretical analysis in terms of cascade smoothing properties.
Finally, concerning the related sense of hearing, briefly touched upon in the beginning of the introduction, let us state that there is a way to define corresponding provably covariant receptive fields for auditory signals, based on the theories in Lindeberg and Friberg (2015) and Lindeberg (2025f), see specifically Section X in the latter reference.